### v1

# Relative periodic solutions in spatial Kepler problem with symmetric perturbation

Abstract.
We study the spatial Kepler problem under a perturbation satisfying both rotational symmetry and reflection symmetry with respect to a plane perpendicular to the rotational axis. By applying recent results from [5] combined with Franks Theorem, we prove the existence of infinitely many relative periodic solutions contained in compact energy surface when the perturbation is sufficiently small and certain technical conditions hold. In addition, we also demonstrate the existence of a unique $z$-symmetric brake orbit which forms a hopf link with a plane relative periodic solutions contained in compact energy surface only requires the perturbation sufficiently small. Our results apply to the motion of a satellite around an ellipsoid with uniform mass distribution and to the $n$-pyramidal problem with one point mass moving along $z$-axis and other $n$ equal masses forming a regular $n$-gon perpendicular to the $z$-axis at all times.
This work is supported by the National Key R&D Program of China (2020YFA0713303). The first and second author is also supported by NSFC (No.12071255). The last author is also supported by NSFC (No. 12171253), Nankai Zhide Fundation and the Fundamental Research Funds for the Central Universities.
1. Introduction
In this paper, we consider the following perturbed spatial Kepler problem
(1) $\ddot{q}=-\frac{q}{|q|^{3}}-\varepsilon\nabla_{q}f(q,\varepsilon),\;q\in\mathbb{R}^{3},$
with the perturbation $f\in C^{\infty}(\mathbb{R}^{3}\backslash\{(0,0,0)\}\times[0,1),\mathbb{R})$ satisfying the following symmetry conditions,
[…]
where $(r,\theta,z)$ is the cylindrical coordinates of $\mathbb{R}^{3}$.
In celestial mechanics, one usually assumes that the celestial bodies are point masses without any size. Such an assumption is justified by the fact that when a celestial body is a perfect sphere with uniform mass distribution, outside the region of such a sphere, its gravitational force is equivalent to a point mass with all the masses concentrated at the center of the sphere.
Meanwhile a planet (or a star) is usually not a perfect sphere, but closer to an ellipsoid with small eccentricity. In this case, the gravitational force can be seen as a Kepler problem with a symmetric perturbation satisfying conditions (2). To be more precise, let’s assume the center of the ellipsoid is located at the origin, and under the spherical coordinates,
$x=\rho\cos\theta\cos\varphi,\;y=\rho\sin\theta\cos\varphi,\;z=\rho\sin\varphi,$
its boundary is a surface of revolution with respect to the $z$-axis satisfying
(3) $\rho=R\left(1-\frac{2}{3}\mathfrak{e}P_{2}(\sin\varphi)\right),\;\text{ where }P_{2}(x)=\frac{3x^{2}-1}{2}.$
Here $R$ denotes the mean radius, $\mathfrak{e}$ the eccentricity, and $P_{2}$ the second Legendre polynomial. For $\mathfrak{e}>0$, the spheroid is oblate (flattened along the $z$-axis), while for $\mathfrak{e}<0$, it is prolate (elongated along the $z$-axis). See [6] for detailed discussions.
For an axially symmetric ellipsoid with total mass $M$ and gravitational constant $G$, the kinetic and potential energies of a massless satellite are given by
(4) $K_{\varpi}(\dot{\rho},\dot{\varphi})$ $=\frac{1}{2}\left(\dot{\rho}^{2}+\rho^{2}\dot{\varphi}^{2}\right)+\frac{\varpi^{2}}{2\rho^{2}\cos^{2}\varphi},$
$U_{\mathfrak{e}}(\rho,\varphi)$ $=-\frac{GM}{\rho}+\frac{2}{5}\mathfrak{e}\frac{GMR^{2}}{\rho^{3}}P_{2}(\sin\varphi)+O(\mathfrak{e}^{2}),$
where $\varpi=\rho^{2}\cos^{2}\varphi\,\dot{\theta}$ represents the conserved angular momentum about the $z$-axis.
Setting $GM=1$ and working in cylindrical coordinates, we observe that the potential function constitutes a small perturbation of the Kepler problem satisfying symmetry conditions (2) when $|\mathfrak{e}|$ is small. We hereafter refer to this system as the ellipsoid problem. For fixed energy $h<0$ and non-zero angular momentum $\varpi$ about the $z$-axis, the energy surface
$\mathcal{M}(h,\varpi,\mathfrak{e})=\left\{(\dot{\rho},\dot{\varphi},\rho,\varphi):K_{\varpi}(\dot{\rho},\dot{\varphi})+U_{\mathfrak{e}}(\rho,\varphi)=h\right\}$
forms a compact manifold homeomorphic to $\mathbb{S}^{3}$. An orbit is called a relative periodic solution if it becomes periodic modulo the $\theta$-coordinate. Our main result states:
Theorem 1.1.
The ellipsoid problem admits infinitely many relative periodic solutions within the compact energy surface $\mathcal{M}(h,\varpi,\mathfrak{e})$ for sufficiently small eccentricity $\mathfrak{e}$.
This result follows as a corollary of our analysis for the perturbed Kepler system (1). To elaborate, we introduce momentum variables associated with cylindrical coordinates $(r,\theta,z)$:
$p_{r}=\dot{r},\quad p_{\theta}=r^{2}\dot{\theta},\quad p_{z}=\dot{z}.$
The corresponding Hamiltonian becomes
(5) $H_{\varepsilon}(p_{r},p_{\theta},p_{z},r,\theta,z)=\frac{1}{2}\left(p_{r}^{2}+r^{-2}p_{\theta}^{2}+p_{z}^{2}\right)-\frac{1}{\sqrt{r^{2}+z^{2}}}+\varepsilon f(r,z,\varepsilon).$
Here $p_{\theta}=r^{2}\dot{\theta}$ represents the angular momentum about the $z$-axis. By fixing this momentum to a nonzero constant $\varpi$, we obtain the reduced Hamiltonian
(6) $H_{\varpi,\varepsilon}(p_{r},p_{z},r,z)=\frac{1}{2}(p_{r}^{2}+p_{z}^{2})+\frac{\varpi^{2}}{2r^{2}}-\frac{1}{\sqrt{r^{2}+z^{2}}}+\varepsilon f(r,z,\varepsilon).$
The corresponding Hamiltonian equation is
(7) $(\dot{p}_{r},\dot{p}_{z},\dot{r},\dot{z})^{T}=J\nabla H_{\varpi,\varepsilon}(p_{r},p_{z},r,z),$
where $J=\begin{pmatrix}0&-I_{2}\\ I_{2}&0\end{pmatrix}$ is the standard symplectic matrix.
We emphasize that solutions of this reduced system correspond to solutions of the original system with fixed angular momentum $\varpi$. In particular, periodic solutions in the reduced system generate relative periodic solutions in the full system through rotational symmetry.
The reduced Hamiltonian system possesses two degrees of freedom, making its energy surface
$\mathcal{M}(h,\varpi,\varepsilon)=\left\{(p_{r},p_{z},r,z):H_{\varpi,\varepsilon}(p_{r},p_{z},r,z)=h\right\}$
a three-dimensional manifold. This configuration exhibits the following fundamental properties,
Proposition 1.1.
For any $\varepsilon\geq 0$ small enough, there is a continuous family of subset $\Gamma_{\varepsilon}\subset\mathbb{R}\setminus\{0\}\times\mathbb{R}$ with
$\Gamma_{0}=\{(\varpi,h)\in\mathbb{R}\setminus\{0\}\times\mathbb{R}:-1<2h\varpi^{2}<0\},$
such that, when $(\varpi,h)\in\Gamma_{\varepsilon}$, the following results hold.
  1. (a)
    $\mathcal{M}(h,\varpi,\varepsilon)$ forms a compact contact manifold homeomorphic to $\mathbb{S}^{3}$ endowed with contact form $\lambda$.
  2. (b)
    There exists $R$ close to $\varpi^{2}/2$ satisfying $\min\Pi_{r}\mathcal{H}(h,\varpi,\varepsilon)>R$, where $\Pi_{r}$ denotes projection to the $r$-axis and
    $\mathcal{H}(h,\varpi,\varepsilon)=\Pi_{r,z}\mathcal{M}(h,\varpi,\varepsilon)=\left\{(r,z):(p_{r},p_{z},r,z)\in\mathcal{M}(h,\varpi,\varepsilon)\right\}.$
Remark $1.$
Here $\mathcal{H}(h,\varpi,\varepsilon)$ represents the Hill’s region. In most of the previous studies of the perturbed Kepler problem, the collision singularity is usually regularized following Levi-Civita [14], Kustaanheimo-Stiefel [13] or Moser [19]. As a result, it is unclear whether the obtained solutions containing collision or not, for examples in [4] and [22]. Meanwhile in our case, the ellipsoid problem is not a point mass. The avoidance of collision is crucial. Property (b) guarantees orbital avoidance of the central body for sufficiently large angular momentum $\varpi$.
This proposition enables the application of symplectic and contact geometric methods to analyze mechanical systems. Recent developments in this direction include notable results for the restricted three-body problem [2, 1, 18] and isosceles three-body problem [9].
The foundational concept of global surfaces of section originated with Poincaré’s work [21] on the circular planar restricted three-body problem. His fundamental theorem on area-preserving twist maps of the annulus, later proved by Birkhoff [3], became known as the Poincaré-Birkhoff theorem. Franks [7] later generalized this result, showing that the existence of one periodic point implies infinitely many periodic points on the annulus. This framework extends to contact manifolds $(\mathcal{M},\lambda)$ homeomorphic to $\mathbb{S}^{3}$: If the Reeb flow admits a disk-like global surface of section, the system must contain either two or infinitely many periodic orbits. Recent breakthrough work by Cristofaro-Gardiner, Hryniewicz, Hutchings, and Liu [5] revealed an algebraic relationship between the contact volume and the action and Seifert rotation number of Reeb orbits (defined in Sections 2 and 3) for contact manifolds with exactly two Reeb orbits. Crucially, violation of this relationship guarantees the existence of infinitely many periodic orbits. In our setting, the reflection symmetry condition in (2) ensures a special periodic solution $\xi_{\varepsilon}(t)$ entirely contained in the $(p_{r},r)$-plane that bounds a global surface of section (see Section 3). Through calculation of this orbit’s action and Seifert rotation number combined with the algebraic relationship, we obtain:
Theorem 1.2.
For sufficiently small $\varepsilon$ and parameters $(\varpi,h)\in\Gamma_{\varepsilon}$, the reduced Hamiltonian system (6) admits infinitely many periodic orbits when one of the conditions from Theorem 3.4 is satisfied.
While the above result proves the existence of infinitely many periodic orbits, it requires some additional conditions. Meanwhile we can show the existence of at least one special periodic orbit, as long as the perturbation is small enough. To explain it, we define two types of periodic solutions as below.
Definition 1.1.
A periodic solution $\xi(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$ of the reduced Hamiltonian system
  1. (a).
    is a brake orbit, if there are two different time moments $t_{1}\neq t_{2}$, such that the momentum variables all vanish, i.e., $p_{r}(t_{i})=p_{z}(t_{i})=0$, $i=1,2$;
  2. (b).
    is a $z$-symmetric orbit, if $p_{r}(t_{0})=z(t_{0})=0$ for some $t_{0}\in\mathbb{R}$ and
    (8) $\xi(t_{0}+t)=\left({\begin{array}[]{cc}A&0\\ 0&-A\\ \end{array}}\right)\xi(t_{0}-t),\quad\forall t\in\mathbb{R},\text{ where }A=\left({\begin{array}[]{cc}-1&0\\ 0&1\\ \end{array}}\right).$
Remark 1.2.
A brake orbit must be a periodic solution and when it is also $z$-symmetric, we say it is a $z$-symmetric brake orbit.
Theorem 1.3.
For sufficiently small $\varepsilon$ and parameters $(\varpi,h)\in\Gamma_{\varepsilon}$, there exists a unique $z$-symmetric brake orbit $\xi_{bz,\varepsilon}(t)$ forming a Hopf link with $\xi_{\varepsilon}(t)$ in $\mathcal{M}(h,\varpi,\varepsilon)$.
Our results extend beyond the ellipsoid problem to a special $(1+n)$-body configuration called the $n$-pyramidal problem. After fixing the center of mass at the origin, there is an invariant sub-system, where one of the masses $m_{0}$ always moving on the $z$-axis and the remaining $n$ equal masses $m$ always forming a regular $n$-gon with the plane containing them always perpendicular to the $z$-axis. Then the movement of one of the mass formed the $n$-gon can represent the movement of whole system. Let $\varepsilon=m/m_{0}$ denote the mass ratio, and fix energy $h<0$ with $z$-axis angular momentum $\varpi\neq 0$ such that the energy surface $\mathcal{M}(h,\varpi,\varepsilon)$ becomes compact, then we have the following Theorem.
Theorem 1.4.
The $n$-pyramidal problem admits infinitely many relative periodic solutions within $\mathcal{M}(h,\varpi,\varepsilon)$ when the mass ratio $\varepsilon\in(0,\varepsilon_{0}(n))$ with sufficiently small $\varepsilon_{0}(n)$ and $2\leq n\leq 472$.
Remark 1.3.
When $n=2$, the $n$-pyramidal problem is the well-known isosceles three body problem, which has been studied by many authors. Moeckel [16] characterized collision asymptotics in this system. Using Maslov-type index (also called Conley-Zehnder index) and contact geometry, the first and third authors and their collaborators [9] established a criterion of existence of infinitely many relative periodic orbits. Grotta-Ragazzo, Liu and Salomão [8] obtained infinitely many relative periodic orbits when the eccentricity of Euler orbit is small enough using Birkhoff normal form. Our theorem solves the problem for eccentricity of Euler orbit $e\in(0,1)$ but mass ratio $\varepsilon$ is small enough.
Remark 1.4.
For techinical reason, the above theorem does not apply when $n>473$. However we believe the same result should still hold.
The paper is organized as follows. In Section 2, we introduce reduced Kepler problem and obtain some results which is needed in the perturbed cases. In Section 3, we consider the perturbed system. Theorem 1.2 and Theorem 1.3 are proved. In Section 4, we introduce the ellipsoid problem and prove Theorem 1.1. In Section 5, we introduce the $n$-pyramidal problem and prove Theorem 1.4.
2. The Kepler Problem
In this section, we obtain some results of the Kepler problem that will be needed in the perturbed cases. Recall that in Cartesian coordinates $q=(x,y,z)\in\mathbb{R}^{3}$ with the corresponding momentum variable $p=(\dot{x},\dot{y},\dot{z})\in\mathbb{R}^{3}$. The Hamiltonian of the Kepler problem can be written as
(9) $H(p,q)=\frac{|p|^{2}}{2}-\frac{1}{|q|}.$
Meanwhile in cylindrical coordinates $(r,\theta,z)$ with $(p_{r},p_{\theta},p_{z})=(\dot{r},r^{2}\dot{\theta},\dot{z})$, the above Hamiltonian can be written correspondingly as
$H(p_{r},p_{\theta},p_{z},r,\theta,z)=\frac{1}{2}(p_{r}^{2}+r^{-2}p_{\theta}^{2}+p_{z}^{2})-\frac{1}{\sqrt{r^{2}+z^{2}}}.$
Notice that $p_{\theta}=r^{2}\dot{\theta}$ is the angular momentum with respect to the $z$-axis, which is a first integral, as the system is symmetric with respect to the $z$-axis. By fixing the angular momentum to be $\varpi$, we get the following reduced Hamiltonian
(10) $H_{\varpi}(p_{r},p_{z},r,z)=\frac{1}{2}(p_{r}^{2}+p_{z}^{2})+\frac{\varpi^{2}}{2r^{2}}-\frac{1}{\sqrt{r^{2}+z^{2}}},$
which is a system with two degrees of freedom. If $h$ is the energy constant, we get the 3 dimensional energy surface
$\mathcal{M}(h,\varpi)=\{(p_{r},p_{z},r,z):H_{\varpi}(p_{r},p_{z},r,z)=h\}.$
Let $\Pi_{r,z}$ denote its projection to the $(r,z)$-plane, we have the Hill’s region as below
$\mathcal{H}(h,\varpi)=\Pi_{r,z}\mathcal{M}(h,\varpi)=\{(r,z):(p_{r},p_{z},r,z)\in\mathcal{M}(h,\varpi)\}.$
Proposition 2.1.
For the reduced Hamiltonian system (10),
  1. (a)
    if $2h\varpi^{2}<-1$, then $\mathcal{M}(h,\varpi)=\varnothing$;
  2. (b)
    if $2h\varpi^{2}=-1$, then $\mathcal{M}(h,\varpi)=\{(0,0,\varpi^{2},0)\}$;
  3. (c)
    if $1<2h\varpi^{2}<0$, then $\mathcal{M}(h,\varpi)$ is a compact manifold homeomorphic to $\mathbb{S}^{3}$ with
    $\min\{r:(p_{r},p_{z},r,z)\in\mathcal{M}(h,\varpi)\}=\frac{\varpi^{2}}{1+\sqrt{1+2h\varpi^{2}}}>\frac{\varpi^{2}}{2};$
  4. (d)
    if $2h\varpi^{2}>0$, then $\mathcal{M}(h,\varpi)$ is unbounded.
Proof.
By a direct computation,
$\nabla H_{\varpi}=\left(p_{r},p_{z},-\frac{\varpi^{2}}{r^{3}}+\frac{r}{(r^{2}+z^{2})^{3/2}},\frac{z}{(r^{2}+z^{2})^{3/2}}\right).$
Then $(0,0,\varpi^{2},0)$ is the unique critical point of $H_{\varpi}$.
Meanwhile the Hessian matrix at $(0,0,\varpi^{2},0)$ is
$D^{2}H_{\varpi}(0,0,\varpi^{2},0)=\left({\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&\frac{1}{\varpi^{6}}&0\\ 0&0&0&\frac{1}{\varpi^{6}}\end{array}}\right),$
which is positive definite. As a result, $\frac{1}{2\varpi^{2}}=H_{\varpi}(0,0,\varpi^{2},0)$ is the minimum of $H_{\varpi}$. Then properties (a) and (b) follow from this immediately.
Notice that $H_{\varpi}$ is increasing with respect to $z^{2}$ and
$\lim_{z\to\infty}H_{\varpi}(p_{r},p_{z},r,z)=\frac{1}{2}(p_{r}^{2}+p_{z}^{2})+\frac{\varpi^{2}}{2r^{2}}.$
Therefor $\mathcal{M}(h,\varpi)$ is bounded in $z$ direction if and only if
$\frac{1}{2\varpi^{2}}<h<\inf\left\{\frac{1}{2}(p_{r}^{2}+p_{z}^{2})+\frac{\varpi^{2}}{2r^{2}}\right\}=0.$
Then $\mathcal{M}(h,\varpi)$ is unbounded, when $2h\varpi^{2}>0$. This proves property (d).
For property (c), we only need to show $\mathcal{M}(h,\varpi)$ is bounded in $r$ direction, when $1<2h\varpi^{2}<0$. Consider the boundary of $\mathcal{H}(h,\varpi)$:
(11) $V_{\varpi}(r,z)=\frac{\varpi^{2}}{2r^{2}}-\frac{1}{\sqrt{r^{2}+z^{2}}}=h.$
Because of $V_{\varpi}(r,z)$ is increasing with respect to $z^{2}$, so we only need to see $z=0$. Solve the equation
(12) $\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}=h,$
we have $r^{\pm}=\frac{\varpi^{2}}{1\pm\sqrt{1+2h\varpi^{2}}}>\frac{\varpi^{2}}{2}$. So we have $\Pi_{r}\mathcal{H}(h,\varpi)=[r^{+},r^{-}]$ is bounded. Therefor $\mathcal{M}(h,\varpi)$ is compact. By Morse theory, $\mathcal{M}(h,\varpi)$ is homeomorphic to $\mathbb{S}^{3}$.
For simplicity, we denote $\mathcal{M}(h,\varpi)$ and $\mathcal{H}(h,\varpi)$ by $\mathcal{M}$ and $\mathcal{H}$ respectively below, when there is no confusion. Without loss of generality, let $\varpi>0$ in the rest of this section.
The Hamiltonian vector field $X_{H_{\varpi}}$ is determined by $\omega(\cdot,X_{H_{\varpi}})=\mathrm{d}H_{\varpi}$, where $\omega=\mathrm{d}p_{r}\wedge\mathrm{d}r+\mathrm{d}p_{z}\wedge\mathrm{d}z$ is the symplectic form. Since $H_{\varpi}$ is a mechanical Hamiltonian, the hypersurface $\mathcal{M}$ is of contact type. That is there exists a 1-form $\lambda$ on $\mathcal{M}$ satisfying $\lambda\wedge\mathrm{d}\lambda\neq 0$, so that its Reeb vector field $R$, determined by $\mathrm{d}(R,\cdot)\equiv 0$ and $\lambda(R)=1$, is parallel to $X_{H_{\varpi}}$ and $\mathrm{d}\lambda=\omega$.
From now on we will focus on the cases with $2h\varpi^{2}\in(-1,0)$. By Proposition 2.1, $\mathcal{M}$ is a compact manifold homeomorphic to $\mathbb{S}^{3}$. Following the definitions in [5], we define its contact volume as
(13) $\text{Vol}(\mathcal{M})=\int_{\mathcal{M}}\lambda\wedge\mathrm{d}\lambda,$
and using Stokes formula
(14) $\text{Vol}(\mathcal{M})=\int_{S}\mathrm{d}\lambda\wedge\mathrm{d}\lambda=2\int_{S}\mathrm{d}p_{r}\wedge\mathrm{d}r\wedge\mathrm{d}p_{z}\wedge\mathrm{d}z,$
where $S$ is the region bounded by $\mathcal{M}$, i.e. $\partial S=\mathcal{M}$.
For any periodic solution $\xi(t)$ of the reduced Kepler problem (10), define its action value as
(15) $A(\xi(t))=\int_{\xi(t)}\lambda.$
For the given energy $h$ and angular momentum $\varpi$, we can find a unique solution $\xi_{0}(t)$ of the reduced Hamiltonian system (10) satisfying
(16) $r(t)=\frac{\varpi^{2}}{1+e\cos\theta(t)},\text{ where }e=\sqrt{1+2h\varpi^{2}}.$
It corresponds to an elliptic orbit of the Kepler problem with eccentrity $e$ and entirely contained in the $(x,y)$-plane. As a result, its action value is
(17) $A(\xi_{0}(t))=\int_{\xi_{0}(t)}\lambda=\int_{\Upsilon}\mathrm{d}\lambda=\int_{\Upsilon}\mathrm{d}p_{r}\wedge\mathrm{d}r,$
where $\Upsilon$ is the region contained in the $(p_{r},r)$-plane and bounded by $\xi_{0}(t)$, i.e. $\partial\Upsilon=\xi_{0}(t)$.
Lemma 2.1.
When $1<2h\varpi^{2}<0$, $\text{Vol}(\mathcal{M})=A^{2}(\xi_{0}(t))$.
Proof.
Notice that $\Upsilon$, $S$ and the Hill’s region $\mathcal{H}$ can be given as
(18) $\Upsilon=\left\{(p_{r},r)|\frac{1}{2}p_{r}^{2}+\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}\leq h\right\},$
$S=\left\{(p_{r},p_{z},r,z)|\frac{1}{2}(p_{r}^{2}+p_{z}^{2})+\frac{\varpi^{2}}{2r^{2}}-\frac{1}{\sqrt{r^{2}+z^{2}}}\leq h\right\},$
$\mathcal{H}=\left\{(r,z)|\frac{\varpi^{2}}{2r^{2}}-\frac{1}{\sqrt{r^{2}+z^{2}}}\leq h\right\}.$
Recall that $r^{\pm}$ are the roots of $\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}=h$. Using (16), we can get
(19) $A(\xi_{0}(t))$ $=\int_{\Upsilon}\mathrm{d}p_{r}\wedge\mathrm{d}r=2\int_{r^{+}}^{r^{-}}\sqrt{2h-\frac{\varpi^{2}}{r^{2}}+\frac{2}{r}}\mathrm{d}r$
$=2\int_{0}^{\pi}\frac{e\sin\theta}{\varpi}\cdot\frac{\varpi^{2}e\sin\theta}{(1+e\cos\theta)^{2}}\mathrm{d}\theta=2e^{2}\varpi\int_{0}^{\pi}\frac{\sin^{2}\theta}{(1+e\cos\theta)^{2}}\mathrm{d}\theta$
$=2\varpi\left(\int_{0}^{\pi}\frac{1}{1+e\cos\theta}\mathrm{d}\theta-\pi\right)=2\pi\varpi\left(\frac{1}{\sqrt{1-e^{2}}}-1\right).$
Meanwhile under polar coordinates $(r,z)=(\rho\cos\varphi,\rho\sin\varphi)$, Hill’s region $\mathcal{H}$ can be given as
$\mathcal{H}=\{(\rho,\varphi)|\rho^{-}(\varphi)\leq\rho\leq\rho^{+}(\varphi),\varphi^{-}\leq\varphi\leq\varphi^{+}\}=\{(\rho,\varphi)|\varphi^{-}(\rho)\leq\varphi\leq\varphi^{+}(\rho),\rho^{-}\leq\rho\leq\rho^{+}\},$
where
$\rho^{\pm}(\varphi)=\frac{1\pm\sqrt{1+2h\varpi^{2}/\cos^{2}\varphi}}{-2h},\qquad\quad\varphi^{\pm}=\pm\arccos\sqrt{-2h\varpi^{2}},$
$\varphi^{\pm}(\rho)=\pm\arctan\sqrt{\frac{2h\rho^{2}+2\rho-\varpi^{2}}{\varpi^{2}}},\quad\rho^{\pm}=\frac{1\pm\sqrt{1+2h\varpi^{2}}}{-2h}.$
Using this we can compute the contact volume of $\mathcal{M}$ as below.
(20) $\text{Vol}(\mathcal{M})$ $=2\int_{S}\mathrm{d}p_{r}\wedge\mathrm{d}r\wedge\mathrm{d}p_{z}\wedge\mathrm{d}z$
$=2\pi\iint_{\mathcal{H}}2h-\frac{\varpi^{2}}{r^{2}}+\frac{2}{\sqrt{r^{2}+z^{2}}}\ \mathrm{d}r\mathrm{d}z$
$=2\pi\iint_{\mathcal{H}}2h\rho-\frac{\varpi^{2}}{\rho\cos^{2}\varphi}+2\ \mathrm{d}\rho\mathrm{d}\varphi$
$=2\pi\left(\int_{\varphi^{-}}^{\varphi^{+}}\int_{\rho^{-}(\varphi)}^{\rho^{+}(\varphi)}2h\rho+2\ \mathrm{d}\rho\mathrm{d}\varphi-\int_{\rho^{-}}^{\rho^{+}}\int_{\varphi^{-}(\rho)}^{\varphi^{+}(\rho)}\frac{\varpi^{2}}{\rho\cos^{2}\varphi}\mathrm{d}\varphi\mathrm{d}\rho\right)$
$=2\pi\left(I_{1}-I_{2}\right)$
To compute $I_{1}$ and $I_{2}$, we need the following formulas, when $a\in(0,1)$,
(21) $\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{\cos^{2}\theta+a\sin^{2}\theta}\mathrm{d}\theta$ $=\int_{-\pi/2}^{\pi/2}\frac{1}{1+a\tan^{2}\theta}\mathrm{d}\theta=\int_{-\infty}^{+\infty}\frac{1}{1+at^{2}}\cdot\frac{1}{1+t^{2}}\mathrm{d}t$
$=\int_{-\infty}^{+\infty}-\frac{a}{1-a}\cdot\frac{1}{1+at^{2}}+\frac{1}{1-a}\cdot\frac{1}{1+t^{2}}\mathrm{d}t$
$=\left(-\frac{\sqrt{a}}{1-a}\arctan\sqrt{a}t+\frac{1}{1-a}\arctan t\right)\bigg{|}^{+\infty}_{-\infty}=\frac{\pi}{1+\sqrt{a}};$
(22) $\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{1+a\sin\theta}\mathrm{d}\theta$ $=\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{1-a^{2}\sin^{2}\theta}\mathrm{d}\theta-\int_{-\pi/2}^{\pi/2}\frac{a\cos^{2}\theta\sin\theta}{1-a^{2}\sin^{2}\theta}\mathrm{d}\theta$
$=\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{\cos^{2}\theta+(1-a^{2})\sin^{2}\theta}\mathrm{d}\theta=\frac{\pi}{1+\sqrt{1-a^{2}}}.$
Since $\cos^{2}\varphi^{+}=-2h\varpi^{2}=1-e^{2}$ and $\sin^{2}\varphi^{+}=1+2h\varpi^{2}=e^{2}$, we choose $\sin\varphi=e\sin\theta$, we conclude that $\cos\varphi\mathrm{d}\varphi=e\cos\theta\mathrm{d}\theta$. we compute that
$I_{1}$ $=\int_{\varphi^{-}}^{\varphi^{+}}\int_{\rho^{-}(\varphi)}^{\rho^{+}(\varphi)}2h\rho+2\mathrm{d}\rho\mathrm{d}\varphi=\frac{1}{-h}\int_{\varphi^{-}}^{\varphi^{+}}\left(1+\frac{2h\varpi^{2}}{\cos^{2}\varphi}\right)^{1/2}\mathrm{d}\varphi$
$=\frac{1}{-h}\int_{\varphi^{-}}^{\varphi^{+}}\frac{(e^{2}-\sin^{2}\varphi)^{1/2}}{\cos^{2}\varphi}\mathrm{d}\sin\varphi=\frac{e^{2}}{-h}\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{1-e^{2}\sin^{2}\theta}\mathrm{d}\theta$
$=\frac{e^{2}}{-h}\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{\cos^{2}\theta+(1-e^{2})\sin^{2}\theta}\mathrm{d}\theta=\frac{e^{2}}{-h}\cdot\frac{\pi}{1+\sqrt{1-e^{2}}}$
$=\frac{2\pi\varpi^{2}}{\sqrt{1-e^{2}}}\cdot\left(\frac{1}{\sqrt{1-e^{2}}}-1\right).$
For $I_{2}$, we choose $u=\rho+1/(2h)=u^{+}\sin\theta$, we have $u^{\pm}=\rho^{\pm}+1/(2h)=\pm e/(-2h)$ and we compute that
$I_{2}$ $=\int_{\rho^{-}}^{\rho^{+}}\int_{\varphi_{-}(\rho)}^{\varphi^{+}(\rho)}\frac{\varpi^{2}}{\rho\cos^{2}\varphi}\mathrm{d}\varphi\mathrm{d}\rho=2\varpi\int_{\rho^{-}}^{\rho^{+}}\frac{\sqrt{2h\rho^{2}+2\rho-\varpi^{2}}}{\rho}\mathrm{d}\rho$
$=2\sqrt{1-e^{2}}\int_{u^{-}}^{u^{+}}\frac{\sqrt{e^{2}/(4h^{2})-u^{2}}}{u+1/(-2h)}\mathrm{d}u=\sqrt{1-e^{2}}\cdot\frac{e^{2}}{-h}\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{1+e\sin\theta}\mathrm{d}\theta$
$=2\pi\varpi^{2}\left(\frac{1}{\sqrt{1-e^{2}}}-1\right).$
The above computation imply $\text{Vol}(\mathcal{M})=A^{2}(\xi_{0}(t))$. ∎
We now study some special orbits, including brake orbits and $z$-symmetric orbits. A definition of a global surface of section is required.
Definition 2.1.
Let $X$ be a non-vanishing vector field on $\mathbb{S}^{3}$ with flow $\varphi_{t}^{X}$. A disk-like global surface of section is an embedded closed 2-dim disk $D\subset\mathbb{S}^{3}$ satisfying
  1. (1)
    $X$ is tangent to $\partial D$, the boundary of $D$;
  2. (2)
    $X$ is transverse to $D^{\circ}$, the interior of $D$;
  3. (3)
    for any $x\in\mathbb{S}^{3}\setminus\partial D$, there exist $t^{+}>0$ and $t^{-}<0$ such that both $\varphi_{t^{+}}^{X}(x)$ and $\varphi_{t^{-}}^{X}(x)$ are contained in the interior of $D$.
Lemma 2.2.
When $2h\varpi^{2}\in(-1,0)$, $\Sigma_{\pm}$ given below, are two disk-like global surfaces of section of the Hamiltonian flow $\varphi_{t}$ corresponding to the Hamiltonian vector field $X_{H_{\varpi}}$,
$\Sigma_{+}=\{(p_{r},p_{z},r,z)\in\mathcal{M}:z=0,p_{z}\geq 0\},$
$\Sigma_{-}=\{(p_{r},p_{z},r,z)\in\mathcal{M}:z=0,p_{z}\leq 0\}.$
Proof.
The boundary $\partial\Sigma_{\pm}=\{z=0,p_{z}=0\}\cap\mathcal{M}(h,\varpi)$ coincides with the orbit $\xi_{0}(t)$, which is tangent to $X_{H_{\varpi}}$.
For $\xi=(p_{r},p_{z},r,z)\in\Sigma_{+}^{\circ}$, the last component of $X_{H_{\varpi}}(\xi)=J\nabla H_{\varpi}(\xi)$ is $p_{z}>0$, ensuring transversality to $\Sigma_{+}^{\circ}$.
Consider a trajectory $\xi(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$ starting from $\Sigma_{+}^{\circ}$. Suppose no $t_{1}>0$ exists with $p_{z}(t_{1})=0$ and $z(t_{1})>0$. Then $z(t)$ monotonically increases to some $z_{0}>0$, and the $\omega$-limit set of $\xi(t)$ lies in $\{z=z_{0}\}\cap\mathcal{M}$. By invariance of $\omega$-limit set under $\varphi_{t}$, there exists a solution $\hat{\xi}(t)=(\hat{p}_{r}(t),\hat{p}_{z}(t),\hat{r}(t),\hat{z}(t))$ with $\hat{z}(t)=z_{0}>0$ and $\hat{p}_{z}(t)=\dot{\hat{z}}(t)=0$. However,
$\dot{\hat{p}}_{z}(t)=\ddot{\hat{z}}(t)=-\frac{\hat{z}(t)}{(\hat{r}^{2}(t)+\hat{z}^{2}(t))^{3/2}}<0,$
yielding a contradiction. Thus, $t_{1}$ exists. Since $\ddot{z}(t)<-cz(t)$ for some $c>0$, by the Sturm-Picone comparison theorem, there exists $t_{+}>t_{1}$ with $p_{z}(t_{+})>0$ and $z(t_{+})=0$, proving $\Sigma_{+}$ is a global surface of section. The argument for $\Sigma_{-}$ is analogous. ∎
Let $\mathcal{B}_{+}=\{(r,z)\in\partial\mathcal{H}:z>0\}$ and $\mathcal{B}_{-}=\{(r,z)\in\partial\mathcal{H}:z<0\}$ denote the upper and lower Hill’s boundaries. For a solution $\xi(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$ with $\Pi_{r,z}\xi(0)\in\mathcal{B}_{+}$, Lemma 2.2 guarantees a minimal $t_{1}>0$ such that $\xi(t_{1})\in\Sigma_{-}$.
Expressing $\xi(t)$ in the original Hamiltonian system (9) via spherical coordinates $(\rho,\theta,\varphi)$, we have
(23) $\rho(t)=\frac{|\Omega|^{2}}{1+\hat{e}\cos{(\hat{\theta}(t)+g)}},$
where $\Omega=q\times\dot{q}=(\varpi_{1},\varpi_{2},\varpi_{3})\in\mathbb{R}^{3}$ is the angular momentum vector, $\hat{e}=\sqrt{1+2h|\Omega|^{2}}$ is the eccentricity, $\hat{\theta}(t)$ satisfies $\rho^{2}(t)\mathrm{d}\hat{\theta}/\mathrm{d}t=|\Omega|$ and $\hat{\theta}(0)=0$ is the true anomaly and $g$ is the argument of perigee. The third component $\varpi=\varpi_{3}$ corresponds to the angular momentum about the $z$-axis.
For $\Pi_{r,z}\xi(0)\in\mathcal{B}_{+}$, the condition $\dot{\rho}(0)=0$ implies $g\in\{0,\pi\}$.
(a) If $g=0$ and $\hat{\theta}(t_{2})=\pi$ for some $t_{2}>t_{1}$, then $\Pi_{r,z}\xi(t_{2})\in\mathcal{B}_{-}$. Here, $\xi(t)$ moves from pericenter to apocenter.
(b) If $g=\pi$ and $\hat{\theta}(t_{2})=\pi$ for some $t_{2}>t_{1}$, then $\Pi_{r,z}\xi(t_{2})\in\mathcal{B}_{-}$. Here, $\xi(t)$ moves from apocenter to pericenter.
In both cases, $p_{z}(t)<0$ for $t\in(0,t_{2})$, and the plane formed by major axis direction of the ellipse orbit and angular momentum direction contains $z$-axis. When $\xi(t)$ intersects the $r$-axis (i.e., at $t=t_{1}$), $\hat{\theta}(t_{1})=\pi/2$. Thus, every trajectory starting from Hill’s boundary is a brake orbit.
For a solution $\bar{\xi}(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$ with $\bar{\xi}(0)\in\Sigma_{+}\cap\{p_{r}=0\}$, Lemma 2.2 ensures a minimal $t_{1}>0$ such that $\bar{\xi}(t_{1})\in\Sigma_{-}$. Since $\bar{\xi}(t)$ is an elliptic orbit, symmetry implies its major axis on the $(x,y)$-plane and $\bar{\xi}(t_{1})\in\Sigma_{-}\cap\{p_{r}=0\}$. Consequently, $\bar{\xi}(t)$ is a $z$-symmetric periodic orbit.
Lemma 2.3.
For a brake orbit $\xi(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$,
$p_{r}(t_{1})=\frac{\varpi^{2}+2hr^{2}(0)}{2\varpi r(0)}.$
Proof.
Interpreting $\xi(t)$ as a solution to the planar Kepler problem,
(24) $H(p_{\rho},p_{\hat{\theta}},\rho,\hat{\theta})=\frac{1}{2}\left(p_{\rho}^{2}+\frac{p_{\hat{\theta}}^{2}}{\rho^{2}}\right)-\frac{1}{\rho},$
and in spherical coordinates $(\rho,\theta,\varphi)$,
(25) $H(p_{\rho},p_{\theta},p_{\varphi},\rho,\theta,\varphi)=\frac{1}{2}\left(p_{\rho}^{2}+\frac{p_{\theta}^{2}}{\rho^{2}\cos^{2}\varphi}+\frac{p_{\varphi}^{2}}{\rho^{2}}\right)-\frac{1}{\rho},$
angular momentum conservation yields
(26) $|\Omega|=|p_{\hat{\theta}}|,\quad\varpi=p_{\theta},$
where $\Omega$ is the total angular momentum vector. Energy conservation implies
(27) $\varpi^{2}\cos^{-2}\varphi+p_{\varphi}^{2}=|\Omega|^{2}.$
[…]
$\frac{\varpi^{2}}{2r^{2}}-\frac{1}{\sqrt{r^{2}+z^{2}}}=h,$
translates to
$z^{2}=r^{2}\left(\left(\frac{2r}{\varpi^{2}-2hr^{2}}\right)^{2}-1\right).$
For $(r,z)\in\mathcal{B}_{+}$, express $r=\rho\cos\varphi$ and $z=\rho\sin\varphi$. Then $\dot{\varphi}=0$, and
$\tan^{2}\varphi=\frac{z^{2}}{r^{2}}=\left(\frac{2r}{\varpi^{2}-2hr^{2}}\right)^{2}-1.$
From (27) and $p_{\varphi}=\rho^{2}\dot{\varphi}=0$, we derive
(28) $\cos^{2}\varphi=\frac{\varpi^{2}}{|\Omega|^{2}},\quad\frac{\varpi}{|\Omega|}=\frac{\varpi^{2}-2hr^{2}}{2r}.$
For $\xi(t_{1})$, also express $\tilde{r}=\tilde{\rho}\cos\tilde{\varphi}$ and $\tilde{z}=\tilde{\rho}\sin\tilde{\varphi}$, $\tilde{z}=0$ implies $\tilde{\varphi}=0$. By (23) and $\cos\hat{\theta}(t_{1})=0$,
(29) $\tilde{r}=|\Omega|^{2},\quad\tilde{p}_{r}=\dot{\tilde{\rho}}=\tilde{p}_{\rho}.$
Energy conservation (24) and angular momentum conservation (26) yield
(30) $\frac{1}{2}\left(\tilde{p}_{r}^{2}+\frac{|\Omega|^{2}}{\tilde{\rho}^{2}}\right)-\frac{1}{\tilde{\rho}}=h.$
Combining (28), (29), and (30), we have
$\tilde{p}_{r}^{2}=\left(\frac{\varpi^{2}+2hr^{2}}{2\varpi r}\right)^{2}.$
To determine the sign of $\tilde{p}_{r}$, observe that near pericenter, $\tilde{\rho}(t)$ increases. Thus, $\tilde{p}_{r}=\tilde{p}_{\rho}>0$, giving
$\tilde{p}_{r}=\frac{\varpi^{2}+2hr^{2}}{2\varpi r}.$
Lemma 2.4.
For $1<2h\varpi^{2}<0$, there exists one and only one $z$-symmetric brake orbit $\xi_{bz}(t)$ besides $\xi_{0}(t)$ inside $\mathcal{M}$, moreover, $\xi_{bz}(t)$ and $\xi_{0}(t)$ forms a hopf link.
Proof.
Assume $\xi(t)$ is a brake orbit. It also a $z$-symmetric orbit only when $p_{r}(t_{1})=0$. By lemma 2.3, we get the unique root $r(0)=\frac{\varpi}{\sqrt{-2h}}$. So there exists one and only one $z$-symmetric brake orbit $\xi_{bz}(t)$ besides $\xi_{0}(t)$ within $\mathcal{M}$. Due to these two orbits linked together exactly once, they formed a hopf link. ∎
Remark 2.1.
The orbit $\xi_{bz}(t)$ corresponds to a circular solution of (9) with angular momentum $\varpi$ about $z$-axis.
3. The Kepler Problem with a symmetric perturbation
Now we will consider the following perturbed Kepler problem
$H_{\varepsilon}(p,q)=\frac{1}{2}|p|^{2}-\frac{1}{|q|}+\varepsilon f(q,\varepsilon).$
with $f$ satisfying conditions (2). Rotational symmetry condition in (2) implies the angular momentum with respect to the $z$-axis is still a first integral. Following the approach given in the previous section, by fixing it to be a non-zero constant $\varpi$, we get a reduced Hamiltonian system
$(\dot{p}_{r},\dot{p}_{z},\dot{r},\dot{z})^{T}=J\nabla H_{\varpi,\varepsilon}(p_{r},p_{z},r,z).$
with the reduced Hamiltonian $H_{\varpi,\varepsilon}$ given as (6).
First, we establish properties analogous to those of the Kepler problem.
Proof of Proposition 1.1.
By Proposition 2.1, when $2h\varpi^{2}\in(-1,0)$, $\mathcal{M}(h,\varpi,0)$ is a compact manifold homeomorphic to $\mathbb{S}^{3}$, where
$\min\left\{r:(p_{r},p_{z},r,z)\in\mathcal{M}(h,\varpi,0)\right\}=\frac{\varpi^{2}}{1+\sqrt{1+2h\varpi^{2}}}>\frac{\varpi^{2}}{2}.$
For $\varepsilon$ sufficiently small, $\mathcal{M}(h,\varpi,\varepsilon)$ remains close to $\mathcal{M}(h,\varpi,0)$.
As in the Kepler problem, the Hamiltonian $H_{\varpi,\varepsilon}$, whose vector field $X_{H_{\varpi,\varepsilon}}$ is defined by $\omega(\ \cdot\ ,X_{H_{\varpi,\varepsilon}})=\mathrm{d}H_{\varpi,\varepsilon}$ with symplectic form $\omega=\mathrm{d}p_{r}\wedge\mathrm{d}r+\mathrm{d}p_{z}\wedge\mathrm{d}z$, is a mechanical Hamiltonian. Consequently, the hypersurface $\mathcal{M}(h,\varpi,\varepsilon)$ is of contact type. Thus, there exists a contact form $\lambda_{\varepsilon}$ on $\mathcal{M}(h,\varpi,\varepsilon)$ satisfying $\lambda_{\varepsilon}\wedge\mathrm{d}\lambda_{\varepsilon}\neq 0$. Its Reeb vector field $R_{\varepsilon}$, determined by $\mathrm{d}\lambda_{\varepsilon}(R_{\varepsilon},\cdot)=0$ and $\lambda_{\varepsilon}(R_{\varepsilon})=1$, is parallel to $X_{H_{\varpi,\varepsilon}}$, with $\mathrm{d}\lambda_{\varepsilon}=\omega$. ∎
Remark 3.1.
While the contact form $\lambda_{\varepsilon}$ and Reeb vector field $R_{\varepsilon}$ differ from those of the Kepler problem, we simplify notation by writing $\lambda$ and $R$ in this section, as computations depend only on the symplectic form $\omega$.
Lemma 3.1.
There always exists a periodic solution $\xi_{\varepsilon}(t)$ entirely contained in the $(p_{r},r)$-plane.
Proof.
Due to the $z$-symmetry of the system, the subsystem restricted to the $(p_{r},r)$-plane is governed by the Hamiltonian
$H_{\varpi,\varepsilon}(p_{r},r)=\frac{1}{2}p_{r}^{2}+\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}+\varepsilon f(r,0,\varepsilon),$
whose solutions is $\xi_{\varepsilon}(t)$. ∎
Lemma 3.2.
The sets
$\Sigma_{+,\varepsilon}$ $=\left\{(p_{r},p_{z},r,z)\in\mathcal{M}(h,\varpi,\varepsilon):z=0,\,p_{z}\geq 0\right\},$
$\Sigma_{-,\varepsilon}$ $=\left\{(p_{r},p_{z},r,z)\in\mathcal{M}(h,\varpi,\varepsilon):z=0,\,p_{z}\leq 0\right\},$
are disk-like global surfaces of section for the reduced Hamiltonian flow $\varphi_{\varepsilon}^{t}$, with $\xi_{\varepsilon}$ as their common boundary.
Proof.
The proof mirrors Lemma 2.2, leveraging the $z$-symmetry and the fact that the system is a perturbation of the Kepler problem. ∎
Recent work by Cristofaro-Gardiner, Hryniewicz, Hutchings, and Liu [5] established an algebraic relationship between the contact volume of a contact manifold and the action and Seifert rotation number of Reeb orbits when the manifold admits exactly two Reeb orbits:
Theorem 3.1.
[5, Theorem 1.2 and 1.5] Let $S$ be a 3-sphere with a contact form $\lambda$ supporting exactly two Reeb orbits, $\xi_{1}$ and $\xi_{2}$. Let $A_{i}$ and $\text{Rot}_{i}$ denote the action and Seifert rotation number of $\xi_{i}$ ($i=1,2$), and let $\text{Vol}(S,\lambda)$ be the contact volume. Then
(31) $\text{Vol}(S,\lambda)=\frac{A_{1}^{2}}{\text{Rot}_{1}}=\frac{A_{2}^{2}}{\text{Rot}_{2}},$
and both Reeb orbits are irrationally elliptic.
Remark 3.2.
We say a Reeb orbit $\xi$ is elliptic if all eigenvalues of its monodromy matrix $\gamma$ lie on $\mathbb{U}$ and semi-simple, where $\mathbb{U}$ denotes the unit circle in the complex plane. If, additionally, the angle $\vartheta$ is irrational for eigenvalues $\sigma=e^{i\vartheta}$ and $\sigma^{-1}=e^{-i\vartheta}$, we say $\xi$ is irrationally elliptic. A Reeb orbit $\xi$ is hyperbolic if all eigenvalues of its monodromy matrix $\gamma$ are real and distinct from $\pm 1$.
By Lemma 3.2, $\Sigma_{-,\varepsilon}$ is a global surface of section. The first return time is defined as
$\hat{\tau}:\Sigma_{-,\varepsilon}\setminus\xi_{\varepsilon}\to\mathbb{R}^{+},\quad z\mapsto\inf\left\{t>0\,\big{|}\,\varphi_{\varepsilon}^{t}(z)\in\Sigma_{-,\varepsilon}\right\},$
and the first return map
$\psi:\Sigma_{-,\varepsilon}\setminus\xi_{\varepsilon}\to\Sigma_{-,\varepsilon}\setminus\xi_{\varepsilon},\quad z\mapsto\varphi_{\varepsilon}^{\hat{\tau}(z)}(z),$
is an area-preserving disk map where periodic points of $\psi$ correspond to periodic orbits of $\varphi_{\varepsilon}^{t}$.
Combined with Franks’ theorem [7]:
Theorem 3.2.
Let $\mathbb{D}$ be an open disk and $\mathfrak{F}:\mathbb{D}\to\mathbb{D}$ an area-preserving disk map. If $\mathfrak{F}$ has two distinct periodic points, then it has infinitely many periodic points.
If the algebraic relation in Theorem 3.1 fails, the system admits infinitely many periodic orbits.
We now analyze how the contact volume of $\mathcal{M}(h,\varpi,\varepsilon)$ and the action/Seifert rotation number of $\xi_{\varepsilon}$ differ from the Kepler problem. Let $\mathcal{M}(h,\varpi,\varepsilon)$ be abbreviated as $\mathcal{M}_{\varepsilon}$.
3.1. Estimate the contact volume and action value
Following the definitions in [5], the contact volume of $\mathcal{M}_{\varepsilon}$ and the action value of $\xi_{\varepsilon}$ are
(32) $\text{Vol}(\mathcal{M}_{\varepsilon})=\int_{\mathcal{M}_{\varepsilon}}\lambda\wedge\mathrm{d}\lambda=\int_{S_{\varepsilon}}\mathrm{d}\lambda\wedge\mathrm{d}\lambda=2\int_{S_{\varepsilon}}\mathrm{d}p_{r}\wedge\mathrm{d}r\wedge\mathrm{d}p_{z}\wedge\mathrm{d}z,$
and
(33) $A(\xi_{\varepsilon}(t))=\int_{\xi_{\varepsilon}(t)}\lambda=\int_{\Upsilon_{\varepsilon}}\mathrm{d}\lambda=\int_{\Upsilon_{\varepsilon}}\mathrm{d}p_{r}\wedge\mathrm{d}r,$
where $S_{\varepsilon}=\{H_{\varpi,\varepsilon}\leq h\}$ and $\Upsilon_{\varepsilon}=S_{\varepsilon}\cap\{p_{z}=0,z=0\}$.
For brevity, let $\text{Vol}(\mathcal{M}_{\varepsilon})$ and $A(\xi_{\varepsilon}(t))$ be abbreviated as $\text{Vol}_{\varepsilon}$ and $A_{\varepsilon}$, respectively.
Lemma 3.3.
When $(\varpi,h)\in\Gamma_{\varepsilon}$, let $r(\theta)=\varpi^{2}/(1+e\cos\theta)$, $e=\sqrt{1+2h\varpi^{2}}$, we have
$\left.\frac{\mathrm{d}\text{Vol}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}=-4\pi\tilde{V}(f),\;\text{ with }\tilde{V}(f)=\iint_{\mathcal{H}}f(r,z,0)\ \mathrm{d}r\mathrm{d}z;$
$\left.\frac{\mathrm{d}A_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}=-\frac{2}{\varpi}\tilde{A}(f),\;\text{ with }\tilde{A}(f)=\int_{0}^{\pi}r^{2}(\theta)f(r(\theta),0,0)\mathrm{d}\theta.$
The proof of Lemma 3.3 relies on the following result:
Lemma 3.4.
Let $F(x,y,\varepsilon)\in C^{\infty}(\mathbb{R}^{2}\times(-\varepsilon_{0},\varepsilon_{0}),\mathbb{R})$ and $D\subset\mathbb{R}^{2}$ is a closed region, $\partial D$ is the boundary of $D$. Assume $\phi^{\varepsilon}$ is a family of differential homeomorphism, $F(x,y,0)|_{\partial D}=0$, and $\left.\frac{\partial F}{\partial x}(x,y,0)\right|_{\partial D}\neq 0$, $\left.\frac{\partial F}{\partial x}(x,y,0)\right|_{\partial D}\neq 0$. then
(34) $\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\int_{\phi^{\varepsilon}D}F(x,y,\varepsilon)\mathrm{d}x\mathrm{d}y=\int_{D}\frac{\partial F}{\partial\varepsilon}(x,y,0)\mathrm{d}x\mathrm{d}y.$
Proof.
Assume $\phi^{\varepsilon}(u,v)=(x(u,v,\varepsilon),y(u,v,\varepsilon))$. Then
(35) $\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\int_{\phi^{\varepsilon}D}F(x,y,\varepsilon)\mathrm{d}x\mathrm{d}y$
$=$ $\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\int_{D}(\phi^{\varepsilon})^{*}F(x,y,0)\mathrm{d}x\wedge\mathrm{d}y+\int_{D}\frac{\partial F}{\partial\varepsilon}(x,y,0)\mathrm{d}x\mathrm{d}y.$
It suffices to show $\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\int_{D}(\phi^{\varepsilon})^{*}F(x,y,0)\mathrm{d}x\wedge\mathrm{d}y=0$. Denote $\left.\frac{\partial x}{\partial\varepsilon}\right|_{\varepsilon=0}$ and $\left.\frac{\partial y}{\partial\varepsilon}\right|_{\varepsilon=0}$ by $s^{1}$ and $s^{2}$ respectively, then
(36) $\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\int_{D}(\phi^{\varepsilon})^{*}F(x,y,0)\mathrm{d}x\wedge\mathrm{d}y$
$=$ $\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\int_{D}F(x(u,v,\varepsilon),y(u,v,\varepsilon),0)\left|\frac{\partial(x,y)}{\partial(u,v)}\right|\mathrm{d}u\mathrm{d}v$
$=$ $\int_{D}\left(\frac{\partial F}{\partial x}s^{1}+\frac{\partial F}{\partial y}s^{2}\right)\left|\frac{\partial(x,y)}{\partial(u,v)}\right|+F\left(\frac{\partial s^{1}}{\partial x}+\frac{\partial s^{2}}{\partial y}\right)\left|\frac{\partial(x,y)}{\partial(u,v)}\right|\mathrm{d}u\mathrm{d}v$
$=$ $\int_{D}\left(\frac{\partial F}{\partial x}s^{1}+F\frac{\partial s^{1}}{\partial x}\right)+\left(\frac{\partial F}{\partial y}s^{2}+F\frac{\partial s^{2}}{\partial y}\right)\mathrm{d}x\mathrm{d}y$
$=$ $\int_{D}\frac{\partial(F\cdot s^{1})}{\partial x}+\frac{\partial(F\cdot s^{2})}{\partial y}\mathrm{d}x\mathrm{d}y$
$=$ $\oint_{\partial D}-F\cdot s^{2}\mathrm{d}x+F\cdot s^{1}\mathrm{d}y.$
Since $F(x,y,0)|_{\partial D}=0$ and $\frac{\partial F}{\partial x},\frac{\partial F}{\partial y}\neq 0$ on $\partial D$, the terms $s^{1}$ and $s^{2}$ remain finite. Thus, the integral vanishes. ∎
Proof of Lemma 3.3.
By (32), we have
(37) $\text{Vol}_{\varepsilon}=2\pi\iint_{\mathcal{H}_{\varepsilon}}2h-\frac{\varpi^{2}}{r^{2}}+\frac{2}{\sqrt{r^{2}+z^{2}}}-2\varepsilon f(r,z,\varepsilon)\ \mathrm{d}r\mathrm{d}z.$
By lemma 3.4, we get
(38) $\left.\frac{\mathrm{d}\text{Vol}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}=-4\pi\iint_{\mathcal{H}}f(r,z,0)\ \mathrm{d}r\mathrm{d}z=-4\pi\tilde{V}(f).$
Here $\mathcal{H}_{\varepsilon}$ is the Hill’s region of $\mathcal{M}_{\varepsilon}$ and $\mathcal{H}$ is the Hill’s region of $\mathcal{M}$.
By (33), then we get
(39) $\left.\frac{\mathrm{d}A_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}$ $=2\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\int_{r_{1,\varepsilon}}^{r_{2,\varepsilon}}\left(2h-\frac{\varpi^{2}}{r^{2}}+\frac{2}{r}-2\varepsilon f(r,0,\varepsilon)\right)^{1/2}\mathrm{d}r$
$=-2\int_{r_{1,0}}^{r_{2,0}}f(r,0,0)\left(2h-\frac{\varpi^{2}}{r^{2}}+\frac{2}{r}\right)^{-1/2}\mathrm{d}r$
$=-\frac{2}{\varpi}\int_{0}^{\pi}r^{2}(\theta)f(r(\theta),0,0)\mathrm{d}\theta=-\frac{2}{\varpi}\tilde{A}(f).$
Here we set $r_{1,\varepsilon},r_{2,\varepsilon}$ are the roots of equation $\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}+\varepsilon f(r,0,\varepsilon)=h$. ∎
3.2. Seifert rotation number of $\xi_{\varepsilon}$
In this subsection, we estimate the Seifert rotation number of $\xi_{\varepsilon}$. Let $\hat{i}(\xi_{\varepsilon})$ denote the mean (Maslov-type) index of the periodic orbit $\xi_{\varepsilon}$. The rotation number of $\xi_{\varepsilon}$ is defined as $\overline{\text{Rot}}(\xi_{\varepsilon})=\frac{\hat{i}(\xi_{\varepsilon})}{2}.$ Since the self-linking number of $\xi_{\varepsilon}$ is 1-1, by Remark 4.4 of [5], the Seifert rotation number of $\xi_{\varepsilon}$ is
(40) $\text{Rot}(\xi_{\varepsilon})=\overline{\text{Rot}}(\xi_{\varepsilon})-1=\frac{\hat{i}(\xi_{\varepsilon})}{2}-1.$
The definitions and properties of the Maslov-type index (also known as the Conley-Zehnder index) and the mean index are listed in Appendix, for more details, we refer the readers to [15]. We denote $\text{Rot}(\xi_{\varepsilon})$ by $\text{Rot}_{\varepsilon}$ below.
To calculate $\text{Rot}_{\varepsilon}$, we analyze the linearized equation along $\xi_{\varepsilon}(t)$
(41) $\dot{y}=JD^{2}H_{\varpi,\varepsilon}(\xi_{\varepsilon}(t))y.$
Using $\theta$ as the time variable, let $T_{\varepsilon}$ denote the period of $\xi_{\varepsilon}(\theta)$. Recall that $r(\theta)=\varpi^{2}/(1+e\cos\theta)$, and let $r_{1,\varepsilon}$ and $r_{2,\varepsilon}$ be the roots of $\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}+\varepsilon f(r,0,\varepsilon)=h.$ From $r^{2}\dot{\theta}=\varpi$, the period $T_{\varepsilon}$ is expressed as
(42) $T_{\varepsilon}$ $=2\int_{r_{1,\varepsilon}}^{r_{2,\varepsilon}}\frac{\varpi}{r^{2}}\cdot\left(2h-\frac{\varpi^{2}}{r^{2}}+\frac{2}{r}-2\varepsilon f(r,0,\varepsilon)\right)^{-1/2}\mathrm{d}r$
$=2\frac{\mathrm{d}}{\mathrm{d}h}\int_{r_{1,\varepsilon}}^{r_{2,\varepsilon}}\frac{\varpi}{r^{2}}\left(2h-\frac{\varpi^{2}}{r^{2}}+\frac{2}{r}-2\varepsilon f(r,0,\varepsilon)\right)^{1/2}\mathrm{d}r,$
yielding
(43) $\tilde{T}(f)=\left.\frac{\mathrm{d}T_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}$ $=-2\frac{\mathrm{d}}{\mathrm{d}h}\int_{r_{1,0}}^{r_{2,0}}f(r,0,0)\frac{\varpi}{r^{2}}\cdot\left(2h-\frac{\varpi^{2}}{r^{2}}+\frac{2}{r}\right)^{-1/2}\mathrm{d}r$
$=-2\frac{\mathrm{d}}{\mathrm{d}h}\int_{0}^{\pi}f(r(\theta),0,0)\mathrm{d}\theta$
$=\frac{2}{e}\int_{0}^{\pi}r^{2}(\theta)\frac{\partial f}{\partial r}(r(\theta),0,0)\cos\theta\mathrm{d}\theta.$
In spherical coordinates $(\rho,\theta,\varphi)$ with momenta $(p_{\rho},p_{\theta},p_{\varphi})=(\dot{\rho},\rho^{2}\cos^{2}\varphi\dot{\theta},\rho^{2}\dot{\varphi})$, the reduced Hamiltonian is
(44) $H_{\varpi,\varepsilon}(p_{\rho},p_{\varphi},\rho,\varphi)=\frac{1}{2}(p_{\rho}^{2}+\frac{p_{\varphi}^{2}}{\rho^{2}})+\frac{\varpi^{2}}{2\rho^{2}\cos^{2}\varphi}-\frac{1}{\rho}+\varepsilon f(\rho,\varphi,\varepsilon).$
Applying the time-scaling transformation $\frac{\mathrm{d}\tau}{\mathrm{d}t}=\frac{\varpi}{\rho^{2}}$, we define
(45) $K_{\varpi,\varepsilon}=\frac{\rho^{2}}{\varpi}(H_{\varpi,\varepsilon}-h)=\frac{1}{\varpi}\left(\frac{1}{2}\rho^{2}p_{\rho}^{2}+\frac{1}{2}{p_{\varphi}^{2}}-h\rho^{2}+\frac{\varpi^{2}}{2\cos^{2}\varphi}-\rho+\varepsilon g(\rho,\varphi,\varepsilon)\right),$
where $g(\rho,\varphi,\varepsilon)=\rho^{2}f(\rho,\varphi,\varepsilon)$. Since
(46) $\nabla K_{\varpi,\varepsilon}=\frac{\rho^{2}}{\varpi}\nabla H_{\varpi,\varepsilon}+\nabla(\rho^{2}/\varpi)(H_{\varpi,\varepsilon}-h),$
the solutions of $\dot{x}=J\nabla H_{\varpi,\varepsilon}(x)$ coincide with those of $\frac{\mathrm{d}x}{\mathrm{d}\tau}=J\nabla K_{\varpi,\varepsilon}(x)$ on the energy surface $K_{\varpi,\varepsilon}^{-1}(0)=H_{\varpi,\varepsilon}^{-1}(h)$. For $\xi_{\varepsilon}$, as a solution of $\frac{\mathrm{d}x}{\mathrm{d}\tau}=J\nabla K_{\varpi,\varepsilon}(x)$, the period in $\tau$ is $T_{\varepsilon}$. The gradient and Hessian of $K_{\varpi,\varepsilon}$ are
$\nabla K_{\varpi,\varepsilon}=\frac{1}{\varpi}\left(\rho^{2}p_{\rho},\,p_{\varphi},\,\rho p_{\rho}^{2}-2h\rho-1+\varepsilon\frac{\partial g}{\partial\rho},\,\frac{\varpi^{2}\sin\varphi}{\cos^{3}\varphi}+\varepsilon\frac{\partial g}{\partial\varphi}\right),$
and
$D^{2}K_{\varpi,\varepsilon}=\frac{1}{\varpi}\begin{pmatrix}\rho^{2}&0&2\rho p_{\rho}&0\\ 0&1&0&0\\ 2\rho p_{\rho}&0&p_{\rho}^{2}-2h+\varepsilon\frac{\partial^{2}g}{\partial\rho^{2}}&\varepsilon\frac{\partial^{2}g}{\partial\rho\partial\varphi}\\ 0&0&\varepsilon\frac{\partial^{2}g}{\partial\rho\partial\varphi}&\frac{\varpi^{2}(1+2\sin^{2}\varphi)}{\cos^{4}\varphi}+\varepsilon\frac{\partial^{2}g}{\partial\varphi^{2}}\end{pmatrix}$.
For $\xi_{\varepsilon}(\tau)$ restricted to the $\varphi$-axis ($\varphi(\xi_{\varepsilon})=0$)
$D^{2}K_{\varpi,\varepsilon}(\xi_{\varepsilon})=\frac{1}{\varpi}\begin{pmatrix}\rho^{2}&0&2\rho p_{\rho}&0\\ 0&1&0&0\\ 2\rho p_{\rho}&0&p_{\rho}^{2}-2h+\varepsilon\frac{\partial^{2}g}{\partial\rho^{2}}&0\\ 0&0&0&\varpi^{2}+\varepsilon\frac{\partial^{2}g}{\partial\varphi^{2}}\end{pmatrix}$.
The linearized equation $\frac{\mathrm{d}y}{\mathrm{d}\tau}=JD^{2}K_{\varpi,\varepsilon}(\xi_{\varepsilon}(\tau))y$ decouples into two independent subsystems
(47) $\frac{\mathrm{d}\eta_{1}}{\mathrm{d}\tau}=\frac{1}{\varpi}J\left({\begin{array}[]{cc}\rho^{2}&2\rho p_{\rho}\\ 2\rho p_{\rho}&p_{\rho}^{2}-2h+\varepsilon\frac{\partial^{2}g}{\partial\rho^{2}}\end{array}}\right)\eta_{1},$
and
(48) $\frac{\mathrm{d}\eta_{2}}{\mathrm{d}\tau}=\frac{1}{\varpi}J\left({\begin{array}[]{cc}1&0\\ 0&\varpi^{2}+\varepsilon\frac{\partial^{2}g}{\partial\varphi^{2}}\end{array}}\right)\eta_{2}.$
Lemma 3.5.
Let $\gamma_{\varepsilon}(\tau)$ be the fundamental solution of (48), then $\text{Rot}_{\varepsilon}=\hat{i}(\gamma_{\varepsilon}(\tau))/2$.
Proof.
Since $\xi_{\varepsilon}(\tau)=(p_{\rho,\varepsilon}(\tau),0,\rho_{\varepsilon}(\tau),0)$ satisfies $\frac{\mathrm{d}\xi_{\varepsilon}(\tau)}{\mathrm{d}\tau}=J\nabla K_{\varpi,\varepsilon}(\xi_{\varepsilon}(\tau))$, differentiating both sides with respect to $\tau$ yields
$\frac{\mathrm{d}^{2}\xi_{\varepsilon}(\tau)}{\mathrm{d}\tau^{2}}=JD^{2}K_{\varpi,\varepsilon}(\xi_{\varepsilon}(\tau))\frac{\mathrm{d}\xi_{\varepsilon}(\tau)}{\mathrm{d}\tau}.$
Thus, $c\cdot\left(\frac{\mathrm{d}p_{\rho,\varepsilon}(\tau)}{\mathrm{d}\tau},\frac{\mathrm{d}\rho_{\varepsilon}(\tau)}{\mathrm{d}\tau}\right)$ is a periodic solution of (47). Because $\frac{\mathrm{d}\rho_{\varepsilon}(\tau)}{\mathrm{d}\tau}$ vanishes twice over a period $T_{\varepsilon}$, equation (47) contributes $2$ to the mean index. By (40), the result follows. ∎
For a $2\times 2$ symplectic path from mechanical systems, the analysis is simpler than the general case. Let $\sigma_{\varepsilon}(\tau)=\rho(\tau)e^{i\vartheta(\tau)}$ and $\sigma_{\varepsilon}^{-1}(\tau)$ denote the eigenvalues of $\gamma_{\varepsilon}(\tau)$, where $\vartheta(t)$ is a continuous non-decreasing function. The mean index is then given by $\hat{i}(\gamma_{\varepsilon}(\tau))=\frac{\vartheta(T_{\varepsilon})-\vartheta(0)}{\pi}.$ However, directly calculating the derivative of rotation number $\text{Rot}_{\varepsilon}$ remains challenging. To address this, we utilize the trace formula introduced by the first author and collaborators [12]. This framework was originally introduced to analyze the linear stability of the Lagrange elliptic relative equilibrium. Below, we briefly summarize relevant results for our purposes. Let $T>0$, $I_{2n}$ denote the $2n\times 2n$ identity matrix, and $J=\begin{pmatrix}0&-I_{n}\\ I_{n}&0\end{pmatrix}$ be the standard symplectic matrix. Let $\mathcal{B}(k)$ denote the space of $k\times k$ symmetric matrices, and let $B(t),D(t)\in C([0,T],\mathcal{B}(2n))$. For the anti-periodic linear Hamiltonian system
(49) $\dot{z}(t)=J(B(t)+\lambda D(t))z(t),\quad z(0)=-z(T),$
the operator $J\frac{\mathrm{d}}{\mathrm{d}t}$ is densely defined in the Hilbert space $E=L^{2}([0,T],\mathbb{C}^{2n})$ with domain
$\mathfrak{D}_{-1}=\left\{z(t)\in W^{1,2}([0,T],\mathbb{C}^{2n})\,\big{|}\,z(0)=-z(T)\right\}.$
Assuming $J\frac{\mathrm{d}}{\mathrm{d}t}-B(t)$ is invertible on $\mathfrak{D}_{-1}$, the operator $\mathcal{F}=D\left(-J\frac{\mathrm{d}}{\mathrm{d}t}-B\right)^{-1}$ is of trace class.
Lemma 3.6.
[12, Theorem 3.3] Let $\gamma_{\lambda}(t)$ is the fundamental solution of (49), then
$\det(id-\lambda\mathcal{F})=\det(\gamma_{\lambda}(T)+I_{2n})\cdot\det(\gamma_{0}(T)+I_{2n})^{-1}.$
Lemma 3.7.
[12, Theorem 2.2] Under the above assumption, for $\lambda$ small enough, we have
$\det(id-\lambda\mathcal{F})=\exp\left\{\sum_{n=1}^{\infty}b_{n}\lambda^{n}\right\},$
where $b_{n}=-\frac{1}{n}Tr(\mathcal{F}^{n})$.
Lemma 3.8.
[12, Corollary 1.3] If $\gamma_{0}(T)=I_{2n}$, then
[…]
where $\mathcal{D}=J\int_{0}^{T}\gamma_{0}^{T}(s)D(s)\gamma_{0}(s)\mathrm{d}s$.
Using the above results, we get the following theorem.
Theorem 3.3.
Define $D(f)$ and $E(f)$ as below
$E(f)=\int_{0}^{\pi}r^{2}(\theta)\frac{\partial^{2}f}{\partial\varphi^{2}}(r(\theta),0,0)\mathrm{d}\theta+\varpi^{2}\tilde{T}(f);$
$D(f)=E^{2}(f)-\left(\int_{0}^{\pi}r^{2}(\theta)\frac{\partial^{2}f}{\partial\varphi^{2}}(r(\theta),0,0)\cos 2\theta\mathrm{d}\theta\right)^{2}.$
Then
$\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}|_{\varepsilon=0}=\begin{cases}0,&\text{ if }D(f)<0;\\ \frac{\text{sign}(E(f))}{2\pi\varpi^{2}}\sqrt{D(f)},&\text{ if }D(f)\geq 0,\end{cases}$
where $\text{sign}(a)=a/|a|$ when $a\neq 0$ and $\text{sign}(0)=0$.
Proof.
To estimate the rotation number $\text{Rot}_{\varepsilon}$ via the trace formula, we reformulate equation (48) as an anti-periodic boundary value problem. First, applying the symplectic transformation $\tilde{\eta}_{2}=\text{diag}(1/\sqrt{\varpi},\sqrt{\varpi})\eta_{2}$, we obtain
$\frac{\mathrm{d}\tilde{\eta}_{2}}{\mathrm{d}\tau}=J\left({\begin{array}[]{cc}1&0\\ 0&1+\frac{\varepsilon}{\varpi^{2}}\cdot\frac{\partial^{2}g}{\partial\varphi^{2}}\end{array}}\right)\tilde{\eta}_{2},$
Reparameterizing $\tau$ to $s$ via $\frac{\mathrm{d}\tau}{\mathrm{d}s}=\frac{T_{\varepsilon}}{2\pi}$ (denoting $^{\prime}=\frac{\mathrm{d}}{\mathrm{d}s}$), the system becomes
(50) $\tilde{\eta}_{2}^{\prime}=\frac{T_{\varepsilon}}{2\pi}J\left({\begin{array}[]{cc}1&0\\ 0&1+\frac{\varepsilon}{\varpi^{2}}\cdot\frac{\partial^{2}g}{\partial\varphi^{2}}\end{array}}\right)\tilde{\eta}_{2},$
where the period of $\xi_{\varepsilon}$ in $s$ is $2\pi$. The fundamental solution of (50) is a reparameterization of $\gamma_{\varepsilon}(\tau)$, denoted $\gamma_{\varepsilon}(s)$ with $s$ as the time variable.
Consider the anti-periodic boundary value problem
(51) $\left\{\begin{aligned} &\tilde{\eta}^{\prime}_{2}=J(I_{2}+\varepsilon D_{\varepsilon})\tilde{\eta}_{2},\\ &\tilde{\eta}_{2}(0)=-\tilde{\eta}_{2}(2\pi).\end{aligned}\right$.
where $D_{\varepsilon}=\text{diag}\left(\frac{\tilde{T}(f)}{2\pi}+O(\varepsilon),\,\frac{\tilde{T}(f)}{2\pi}+\frac{T_{\varepsilon}}{2\varpi^{2}\pi}\cdot\frac{\partial^{2}g}{\partial\varphi^{2}}+O(\varepsilon)\right)$ and $\tilde{T}(f)=\left.\frac{\mathrm{d}T_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}$. Since the operator $J\frac{\mathrm{d}}{\mathrm{d}s}-I_{2}$ is invertible in $\mathfrak{D}_{-1}$, Lemma 3.6 implies that for $\mathcal{F}_{\varepsilon}=D_{\varepsilon}\left(-J\frac{\mathrm{d}}{\mathrm{d}s}-I_{2}\right)^{-1}$,
$\det(id-\varepsilon\mathcal{F}_{\varepsilon})=\det(\gamma_{\varepsilon}(2\pi)+I_{2})\cdot\det(\gamma_{0}(2\pi)+I_{2})^{-1}.$
Direct computation shows $\gamma_{0}(s)=\begin{pmatrix}\cos s&-\sin s\\ \sin s&\cos s\end{pmatrix}$ and $\gamma_{0}(2\pi)=I_{2}$. Let $\sigma_{\varepsilon}$ and $\sigma^{-1}_{\varepsilon}$ denote the eigenvalues of $\gamma_{\varepsilon}(2\pi)$. By Lemmas 3.7 and 3.8,
$1-\frac{1}{2}Tr(\mathcal{F}_{\varepsilon}^{2})\varepsilon^{2}+o(\varepsilon^{2})=\frac{2+\sigma_{\varepsilon}+\sigma_{\varepsilon}^{-1}}{4}.$
Differentiating twice with respect to $\varepsilon$ at $\varepsilon=0$,
$Tr(\mathcal{F}_{0}^{2})=(\sigma_{0}^{\prime})^{2}/2.$
Using Lemma 3.8 and the fact that $r(\theta)=\varpi^{2}/(1+e\cos\theta)$ is an even $2\pi$-periodic function, we compute
$\left(\frac{1}{\varpi^{2}}\int_{0}^{\pi}r^{2}(\theta)\frac{\partial^{2}f}{\partial\varphi^{2}}(r(\theta),0,0)\mathrm{d}\theta+\tilde{T}(f)\right)^{2}+\frac{1}{\varpi^{4}}\left(\int_{0}^{\pi}r^{2}(\theta)\frac{\partial^{2}f}{\partial\varphi^{2}}(r(\theta),0,0)\cos 2\theta\mathrm{d}\theta\right)^{2}=(\sigma_{0}^{\prime})^{2}.$
i.e.,
(52) $D(f)/\varpi^{4}=(\sigma_{0}^{\prime})^{2}.$
Case 1: $D(f)<0$. Then $\sigma_{0}^{\prime}$ is real, all eigenvalues of $\gamma_{\varepsilon}(2\pi)$ are real, and $\text{Rot}_{\varepsilon}=1=\text{Rot}_{0}$ with $\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}=0$.
Case 2: $D(f)=0$. Then $\sigma_{0}^{\prime}=0$, and the eigenvalue variation of $\gamma_{\varepsilon}(2\pi)$ is of higher order in $\varepsilon$, so $\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}=0$.
Case 3: $D(f)>0$. Then $\sigma_{0}^{\prime}$ is purely imaginary. Writing $\sigma_{\varepsilon}=\exp(ic_{\varepsilon})$, we have $(\sigma_{0}^{\prime})^{2}=-(c_{0}^{\prime})^{2}$, giving $\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}=\pm\frac{|c_{0}^{\prime}|}{2\pi}=\pm\frac{1}{2\pi\varpi^{2}}\sqrt{D(f)}$. To determine the sign, rewrite (51) as
(53) $\left\{\begin{aligned} &\tilde{\eta}_{2}^{\prime}=J\left(I_{2}+\varepsilon\frac{E(f)}{2\pi\varpi^{2}}+\varepsilon\left(D_{\varepsilon}-\frac{E(f)}{2\pi\varpi^{2}}\right)\right)\tilde{\eta}_{2},\\ &\tilde{\eta}_{2}(0)=-\tilde{\eta}_{2}(2\pi).\end{aligned}\right$.
and analyze the modified system
(54) $\left\{\begin{aligned} &\tilde{\eta}_{2}^{\prime}=J\left(I_{2}+\varepsilon\left(D_{\varepsilon}-\frac{E(f)}{2\pi\varpi^{2}}\right)\right)\tilde{\eta}_{2},\\ &\tilde{\eta}_{2}(0)=-\tilde{\eta}_{2}(2\pi).\end{aligned}\right$.
Repeating the eigenvalue analysis for (54), the monodromy matrix eigenvalues are real. By Sturm’s comparison theorem: For $E(f)>0$, $\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}>0$. For $E(f)<0$, $\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}<0$. ∎
Remark $3.$
Although D(f)=D(f)D(f)=D(-f), the derivative of $\text{Rot}_{\varepsilon}$ is opposite.
The above result implies the following corollary.
Corollary 3.1.
$\xi_{\varepsilon}(t)$ is elliptic, when $D(f)>0$, and hyperbolic, when $D(f)<0$.
Remark 3.4.
We say $\xi_{\varepsilon}(t)$ is linear stable, if $\xi_{\varepsilon}(t)$ is elliptic, and linear unstable, if $\xi_{\varepsilon}(t)$ is hyperbolic.
3.3. The existence of infinitely many periodic orbits and a unique brake $z$-symmetric orbit which forms a hopf link with $\xi_{\varepsilon}$.
Now we will show the perturbed system has infinitely many periodic solutions under certain conditions based on Theorem 3.1 and Theorem 3.2.
Theorem 3.4.
For sufficiently small $\varepsilon$ and parameters $(\varpi,h)\in\Gamma_{\varepsilon}$, let $C(e)=\frac{1}{\sqrt{1-e^{2}}}-1$, the reduced Hamiltonian system (7) admits infinitely many periodic orbits within $\mathcal{M}_{\varepsilon}$, if one of the following two conditions holds
  1. (i)
    $D(f)<0$;
  2. (ii)
    $D(f)\geq 0$ and $\tilde{V}(f)\neq 2C(e)\tilde{A}(f)+C^{2}(e)\text{sign}(E(f))\sqrt{D(f)}/2$.
Proof.
First, from Lemma 2.1 and the fact that $\text{Rot}_{0}=1$, we conclude that the Kepler problem satisfies the algebraic relations stated in Theorem 3.1. For the perturbed system (7), by Theorems 3.1 and 3.2, if the derivatives of both sides with respect to $\varepsilon$ are unequal, i.e.
$\left.\frac{\mathrm{d}\text{Vol}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\neq 2A(\xi_{r}(t))\left.\frac{\mathrm{d}A_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}-A^{2}(\xi_{r}(t))\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0},$
then infinitely many periodic orbits exist for sufficiently small $\varepsilon$. By Theorem 3.3, Lemma 3.3, and equation (19), this inequality is equivalent to conditions (ii). Furthermore, by Corollary 3.1, Theorem 3.1, and Theorem 3.2, condition (i) implies that $\xi_{\varepsilon}$ is hyperbolic and that infinitely many periodic orbits exist. ∎
Remark 3.5.
From previous definition, $\tilde{V}(f)$, $\tilde{A}(f)$ and $\sqrt{D(f)}$ are the scaled derivative of $\text{Vol}_{\varepsilon}$, $A_{\varepsilon}$ and $\text{Rot}_{\varepsilon}$ correspondingly.
While one can obtain infinitely many periodic orbits with the above theorem, it requires some additional conditions besides the perturbation is small enough. Meanwhile our next result shows the existence of at least one periodic orbit when the symmetric perturbation is small enough without any further assumption.
Proof of Theorem 1.3.
We apply the Implicit Function Theorem (IFT) to prove this theorem. First, we analytically construct a one-to-one correspondence between the $r$-coordinates associated with $\mathcal{B}_{+,\varepsilon}$ and $\mathcal{B}_{+,0}$, where
$\mathcal{B}_{+,\varepsilon}=\left\{\frac{\varpi^{2}}{2r^{2}}-\frac{1}{\sqrt{r^{2}+z^{2}}}+\varepsilon f(r,z,\varepsilon)=h\right\}$
is the upper boundary of the Hill’s region $\mathcal{H}_{\varepsilon}$, while $\mathcal{B}_{+,0}$ denotes the upper boundary of the Hill’s region $\mathcal{H}$ for the unperturbed Kepler problem. Among possible approaches to establish this correspondence, we parametrize points on $\mathcal{B}_{+,\varepsilon}$ and $\mathcal{B}_{+,0}$ by constructing rays from $(\varpi^{2},0)$ that intersect these boundaries at $(R(r,\varepsilon),Z(r,\varepsilon))$ and $(r,z(r))$, respectively.
[…]
For every $(r,z(r))\in\mathcal{B}_{+,0}$, we have $R(r,0)=r$ and $Z(r,0)=z(r)$. Define the function
$f_{1}(t,r,\varepsilon)=\Pi_{z}\varphi^{t}_{\varepsilon}(0,0,R(r,\varepsilon),Z(r,\varepsilon)).$
When $\varepsilon=0$, Lemma 2.2 ensures that $\Sigma_{-}$ is a global surface of section. Hence, for fixed $r$, there exists a smallest $T(r)>0$ such that
$\varphi_{0}^{T(r)}(0,0,r,z(r))\in\Sigma_{-}^{\circ}.$
This implies:
$f_{1}(T(r),r,0)=\Pi_{z}\varphi_{0}^{T(r)}(0,0,r,z(r))=0,$
$\partial_{t}f_{1}(T(r),r,0)=\Pi_{p_{z}}\varphi_{0}^{T(r)}(0,0,r,z(r))<0.$
By the IFT, there exists a unique $\tilde{T}(r,\varepsilon)$ near $(r,0,T(r))$ such that $\tilde{T}(r,0)=T(r)$ and
$\Pi_{z}\varphi_{\varepsilon}^{\tilde{T}(r,\varepsilon)}(0,0,R(r,\varepsilon),Z(r,\varepsilon))=0.$
Next, consider the function
$f_{2}(r,\varepsilon)=\Pi_{p_{r}}\varphi_{\varepsilon}^{\tilde{T}(r,\varepsilon)}(0,0,R(r,\varepsilon),Z(r,\varepsilon)).$
For $\varepsilon=0$, let $r_{1}=\frac{\varpi}{\sqrt{-2h}}$. By Lemmas 2.3 and 2.4,
$f_{2}(r_{1},0)=\Pi_{p_{r}}\varphi_{0}^{T(r_{1})}(0,0,r_{1},z(r_{1}))=\frac{\varpi^{2}+2hr_{1}^{2}}{2\varpi r_{1}}=0,$
$\partial_{r}f_{2}(r_{1},0)=\frac{2h}{\varpi}\neq 0.$
Applying the IFT again, there exists a unique $\tilde{r}(\varepsilon)$ near $(0,r_{1})$ such that $\tilde{r}(0)=r_{1}$ and
$\Pi_{p_{r}}\varphi_{\varepsilon}^{\tilde{T}(\tilde{r}(\varepsilon),\varepsilon)}(0,0,R(\tilde{r}(\varepsilon),\varepsilon),Z(\tilde{r}(\varepsilon),\varepsilon))=0.$
This guarantees the existence of a unique $z$-symmetric brake orbit $\xi_{bz,\varepsilon}(t)$ forming a Hopf link with $\xi_{\varepsilon}(t)$. ∎
Remark 3.6.
There may be many even infinitely many $z$-symmetric brake orbits, but only $\xi_{bz,\varepsilon}(t)$ forms a Hopf link with $\xi_{\varepsilon}(t)$.
4. The ellipsoid problem
In this section, we apply results from the previous section to the ellipsoid problem. Recall that when the eccentricity $\mathfrak{e}$ is samll enough, the potential function can be written as (4). If we further assume $GM=1$, then
(55) $U(\rho,\varphi)$ $=-\frac{1}{\rho}+\frac{R^{2}}{5}\mathfrak{e}\frac{3\sin^{2}\varphi-1}{\rho^{3}}+O(\mathfrak{e}^{2})$
$=-\frac{1}{\rho}+\frac{3\sin^{2}\varphi-1}{\rho^{3}}\varepsilon+O(\varepsilon^{2}).$
where $\varepsilon=\frac{R^{2}}{5}\mathfrak{e}$ as a small parameter.
Theorem 1.1 follows from the next result.
Theorem 4.1.
For sufficiently small $\varepsilon$, when $(\varpi,h)\in\Gamma_{\varepsilon}$,
  1. (a)
    the ellipsoid problem have a solution $\xi_{\varepsilon}(t)$ which locate on $r$-axis and it is linear stable.
  2. (b)
    there exists one and only one $z$-symmetric brake orbit $\xi_{bz,\varepsilon}(t)$ in $\mathcal{M}_{\varepsilon}$ which forms a hopf link with $\xi_{\varepsilon}(t)$.
  3. (c)
    the ellipsoid problem admits infinitely many relative periodic solutions within $\mathcal{M}_{\varepsilon}$.
Proof.
Interpreting this system as a Kepler problem with symmetric perturbation, we express the perturbation term as
$f(r,z,\varepsilon)=\frac{2z^{2}-r^{2}}{(r^{2}+z^{2})^{5/2}}+O(\varepsilon)=\frac{3\sin^{2}\varphi-1}{\rho^{3}}+O(\varepsilon).$
The existence of $\xi_{\varepsilon}$ follows from Lemma 3.1. Then Theorem 1.3 implies property (b). To establish property (c), we apply Theorem 3.4 through the following calculations.
First, determine the sign of $D(f)$ by computing
$\tilde{T}(f)=\frac{2}{e}\int_{0}^{\pi}\frac{3}{r^{2}(\theta)}\cos\theta\mathrm{d}\theta=\frac{6\pi}{\varpi^{4}},\quad E(f)=\int_{0}^{\pi}\frac{6}{r(\theta)}\mathrm{d}\theta+\frac{6\pi}{\varpi^{2}}=\frac{12\pi}{\varpi^{2}},$
we obtain
$D(f)=\left(\frac{12\pi}{\varpi^{2}}\right)^{2}-\left(\int_{0}^{\pi}\frac{6}{r(\theta)}\cos 2\theta\mathrm{d}\theta\right)^{2}=\left(\frac{12\pi}{\varpi^{2}}\right)^{2}>0.$
By Corollary 3.1, $\xi_{\varepsilon}(t)$ is linear stable and this proves property (a). Then we need to verify the condition (ii) in Theorem 3.4.
We compute $\tilde{V}(f)$ using polar coordinates with the substitution $\sin\varphi=e\sin\theta$ where $e=\sqrt{1+2h\varpi^{2}}$,
(56) $\tilde{V}(f)$ $=\int_{\mathcal{H}}f(r,z,0)\mathrm{d}r\mathrm{d}z=\int_{\mathcal{H}}\frac{3\sin^{2}\varphi-1}{\rho^{2}}\mathrm{d}\rho\mathrm{d}\varphi$
$=\frac{2}{\varpi^{2}}\int_{\varphi^{-}}^{\varphi^{+}}(3\sin^{2}\varphi-1)\cos^{2}\varphi\left(1+\frac{2h\varpi^{2}}{\cos^{2}\varphi}\right)^{1/2}\mathrm{d}\varphi$
$=\frac{2e^{2}}{\varpi^{2}}\int_{-\pi/2}^{\pi/2}(3e^{2}\sin^{2}\theta-1)\cos^{2}\theta\mathrm{d}\theta$
$=-\frac{\pi}{4\varpi^{2}}e^{2}(4-3e^{2}).$
Then we study the right side,
$2C(e)\tilde{A}(f)+\frac{1}{2}C^{2}(e)\sqrt{D(f)}=\frac{2\pi C(e)}{\varpi^{2}}(3C(e)-1),$
with $\tilde{A}(f)=-\int_{0}^{\pi}\frac{1}{r(\theta)}\mathrm{d}\theta=-\frac{\pi}{\varpi^{2}}$. Crucially, the inequality
(57) $\frac{\pi}{4\varpi^{2}}e^{2}(4-3e^{2})<\frac{2\pi C(e)}{\varpi^{2}}(3C(e)-1)\quad\forall e\in(0,1),$
implies condition (c) holds. ∎
5. The n-pyramidal problem
In this section, we apply the preceding results to the $n$-pyramidal problem. Recall that the fixed axis is $z$-axis and the masses of the $1+n$ bodies are $m_{0},m_{1},\cdots,m_{n}$ and $m_{j}=m$ for $j=1,2,\cdots,n$. Let $q_{1}=(x,y,z)\in\mathbb{R}^{3}$ is the position of $m_{1}$, since the mass center is fixed at the origin, the position of $m_{0}$ is $q_{0}=(0,0,-n\varepsilon z)$, where $\varepsilon=m/m_{0}$ is the mass ratio. Correspondingly, the velocity of $m_{1}$: $\dot{q}_{1}=(\dot{x},\dot{y},\dot{z})$ and $m_{0}$: $\dot{q}_{0}=(0,0,-n\varepsilon\dot{z})$. Then the system can be described by the motion of $m_{1}$ and the Lagrangian is
(58) $\hat{L}(q,\dot{q})$ $=\frac{nm}{2}(\dot{x}^{2}+\dot{y}^{2}+(1+n\varepsilon)\dot{z}^{2})$
$+\frac{1}{2}m^{2}\left(\frac{nM(n)}{\sqrt{x^{2}+y^{2}}}+\frac{2n\varepsilon^{-1}}{\sqrt{x^{2}+y^{2}+(1+n\varepsilon)^{2}z^{2}}}\right),$
where
(59) $M(n)=\frac{1}{2}\sum\limits_{i=1}\limits^{n-1}\csc(\frac{i\pi}{n}).$
Assuming $m_{0}=1$ without loss of generality (implying $m=\varepsilon$), we define the normalized Lagrangian $\tilde{L}=\hat{L}/(nm)$. Under cylindrical coordinates
$x=r\cos\theta,\quad y=r\sin\theta,\quad z=z,$
this becomes
$\tilde{L}(r,\theta,z,\dot{r},\dot{\theta},\dot{z})=\frac{1}{2}(\dot{r}^{2}+r^{2}\dot{\theta}^{2}+(1+n\varepsilon)\dot{z}^{2})+\frac{M(n)\varepsilon}{2r}+\frac{1}{\sqrt{r^{2}+(1+n\varepsilon)^{2}z^{2}}}.$
Through the Legendre transformation
$p_{r}=\dot{r},\quad p_{\theta}=r^{2}\dot{\theta},\quad p_{z}=\dot{z},$
with scaled coordinate $\hat{z}=\sqrt{1+n\varepsilon}\,z$ and fixed angular momentum $r^{2}\dot{\theta}=\varpi$, we obtain the reduced Hamiltonian,
(60) $H_{\varpi,\varepsilon}(p_{r},p_{\hat{z}},r,\hat{z})=\frac{1}{2}(p_{r}^{2}+p_{\hat{z}}^{2})+\frac{\varpi^{2}}{2r^{2}}-\frac{M(n)\varepsilon}{2r}-\frac{1}{\sqrt{r^{2}+(1+n\varepsilon)\hat{z}^{2}}}.$
Hereafter, we denote $\hat{z}$ simply by $z$ when no confusion arises. The parameter set $\Gamma_{\varepsilon}$ can be explicitly calculated for this system.
Proposition 5.1.
For $(1+(M(n)\varepsilon)/2)^{2}<2h\varpi^{2}<-(M(n)\varepsilon)^{2}/4$, energy surface $\mathcal{M}_{\varepsilon}$ is a compact manifold homeomorphic to $\mathbb{S}^{3}$ with contact form $\lambda$.
Proof.
The proof is similar to Proposition 2.1, leaving it to the readers. ∎
Theorem 5.1.
For sufficiently small $\varepsilon$ and $2h\varpi^{2}\in(-\frac{(2+M(n)\varepsilon)^{2}}{4},-\frac{(M(n)\varepsilon)^{2}}{4})$ with $2\leq n\leq 472$,
  1. (a)
    the n-pyramidal problem have a $\varpi$ relative periodic solution $\xi_{\varepsilon}(t)$ which locate on $r$-axis and it is linear stable.
  2. (b)
    there exists one and only one $z$-symmetric brake orbit $\xi_{bz,\varepsilon}(t)$ in $\mathcal{M}_{\varepsilon}$ which forms a hopf link with $\xi_{\varepsilon}(t)$.
  3. (c)
    the n-pyramidal problem admits infinitely many $\varpi$ relative periodic solutions in $\mathcal{M}_{\varepsilon}$.
This result clearly implies Theorem 1.4.
Remark 5.1.
The linear stability of $\xi_{\varepsilon}(t)$ is refer to the linear stability in this sub-manifold. The global linear stability of this system can see [17],[10],[11] and [20].
Proof.
We interpret this system as a Kepler problem with symmetric perturbation, where the perturbation term can be written as
$f(r,z,\varepsilon)=f(\rho,\varphi,\varepsilon)=-\frac{M(n)}{2\rho\cos\varphi}+\frac{n\sin^{2}\varphi}{2\rho}+O(\varepsilon),$
$f(r,0,0)=-\frac{M(n)}{2r}.$
The existence of $\xi_{\varepsilon}$ follows from Lemma 3.1. Then Theorem 1.3 implies property (b). To establish property (c), we apply Theorem 3.4 through the following analysis.
First, we determine the sign of $D(f)$ by computing
$\tilde{T}(f)=\frac{2}{e}\int_{0}^{\pi}r^{2}(\theta)\cdot\frac{M(n)}{2r^{2}(\theta)}\cos\theta\mathrm{d}\theta=0,\quad E(f)=\left(n-\frac{M(n)}{2}\right)\left(\int_{0}^{\pi}r(\theta)\mathrm{d}\theta\right),$
we obtain
$D(f)$ $=\left(n-\frac{M(n)}{2}\right)^{2}\left(\int_{0}^{\pi}r(\theta)\mathrm{d}\theta\right)^{2}-\left(n-\frac{M(n)}{2}\right)^{2}\left(\int_{0}^{\pi}r(\theta)\cos 2\theta\mathrm{d}\theta\right)^{2}$
$=\left(n-\frac{M(n)}{2}\right)^{2}\cdot\frac{\pi^{2}\varpi^{4}}{1-e^{2}}\left(1-\left(\frac{-e}{1+\sqrt{1-e^{2}}}\right)^{4}\right)$
$=\frac{(2n-M(n))^{2}\pi^{2}\varpi^{4}G^{2}(e)}{4(1-e^{2})}>0,$
where $G(e)=\sqrt{1-\left(\frac{-e}{1+\sqrt{1-e^{2}}}\right)^{4}}$. Here we used the fact that $\int_{0}^{\pi}\frac{\cos nt}{1+e\cos t}dt=\frac{\pi}{\sqrt{1-e^{2}}}\left(\frac{-e}{1+\sqrt{1-e^{2}}}\right)^{n}$ for $e\in(0,1)$. By Corollary 3.1, $\xi_{\varepsilon}(t)$ is linear stable and this implies property (a).
When $2\leq n\leq 472$, this equivalent to $2n>M(n)$, therefor $E(f)>0$ and we need verify condition (ii) in Theorem 3.4. Then we calculate $\tilde{V}(f)$,
$\tilde{V}(f)=\int_{\mathcal{H}}f(r,z,0)\mathrm{d}r\mathrm{d}z=-\frac{M(n)}{2}\int_{\mathcal{H}}\frac{1}{\cos\varphi}\mathrm{d}\rho\mathrm{d}\varphi+\frac{n}{2}\int_{\mathcal{H}}\sin^{2}\varphi\mathrm{d}\rho\mathrm{d}\varphi.$
For $\int_{\mathcal{H}}\sin^{2}\varphi\mathrm{d}\rho\mathrm{d}\varphi$, using variable substitution $\sin\varphi=e\sin\theta$ and formula (21),
$\int_{\mathcal{H}}\sin^{2}\varphi\mathrm{d}\rho\mathrm{d}\varphi=\frac{1}{-h}\int_{\varphi^{-}}^{\varphi^{+}}\sin^{2}\varphi\left(1+\frac{2h\varpi^{2}}{\cos^{2}\varphi}\right)^{1/2}\mathrm{d}\varphi$
$=$ $\frac{1}{-h}\int_{\varphi^{-}}^{\varphi^{+}}\tan^{2}\varphi(e^{2}-\sin^{2}\varphi)^{1/2}\mathrm{d}\sin\varphi$
$=$ $\frac{1}{-h}\int_{\varphi^{-}}^{\varphi^{+}}(\sec^{2}\varphi-1)(e^{2}-\sin^{2}\varphi)^{1/2}\mathrm{d}\sin\varphi$
$=$ $\frac{e^{2}}{-h}\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{\cos^{2}\theta+(1-e^{2})\sin^{2}\theta}-\cos^{2}\theta\mathrm{d}\theta$
$=$ $\frac{e^{2}}{-h}\left(\frac{\pi}{1+\sqrt{1-e^{2}}}-\frac{\pi}{2}\right)=\pi\varpi^{2}C^{2}(e).$
For $\int_{\mathcal{H}}\frac{1}{\cos\varphi}\mathrm{d}\rho\mathrm{d}\varphi$, using the inequality $(1+\tan^{2}\varphi)^{3/2}\geq(1+\frac{3}{2}\tan^{2}\varphi)$,
$\int_{\mathcal{H}}\frac{1}{\cos\varphi}\mathrm{d}\rho\mathrm{d}\varphi=\frac{1}{-h}\int_{\varphi^{-}}^{\varphi^{+}}\frac{1}{\cos\varphi}\left(1+\frac{2h\varpi^{2}}{\cos^{2}\varphi}\right)^{1/2}\mathrm{d}\varphi$
$=$ $\frac{1}{-h}\int_{\varphi^{-}}^{\varphi^{+}}(1+\tan^{2}\varphi)^{3/2}(e^{2}-\sin^{2}\varphi)^{1/2}\mathrm{d}\sin\varphi$
$\geq$ $\frac{1}{-h}\int_{\varphi^{-}}^{\varphi^{+}}(1+\frac{3}{2}\tan^{2}\varphi)(e^{2}-\sin^{2}\varphi)^{1/2}\mathrm{d}\sin\varphi$
$=$ $\pi\varpi^{2}C(e)\left(\frac{1}{\sqrt{1-e^{2}}}+1\right)+\frac{3}{2}\pi\varpi^{2}C^{2}(e)$
$=$ $\frac{\pi}{2}\varpi^{2}C^{2}(e)+\frac{2\pi\varpi^{2}C(e)}{\sqrt{1-e^{2}}}.$
Combine above computations, we obtain
$\tilde{V}(f)\leq\frac{\pi}{4}(2n-M(n))\varpi^{2}C^{2}(e)-\frac{M(n)\pi\varpi^{2}C(e)}{\sqrt{1-e^{2}}}.$
Then we calculate the right side,
$\tilde{A}(f)=\int_{0}^{\pi}r^{2}(\theta)f(r(\theta),0,0)\mathrm{d}\theta=-\frac{M(n)}{2}\int_{0}^{\pi}r(\theta)\mathrm{d}\theta=-\frac{M(n)\pi\varpi^{2}}{2\sqrt{1-e^{2}}},$
and
$2C(e)\tilde{A}(f)+C^{2}(e)\sqrt{D(f)}/2\\ =\frac{\pi}{4}(2n-M(n))\varpi^{2}C^{2}(e)\frac{G(e)}{\sqrt{1-e^{2}}}-\frac{M(n)\pi\varpi^{2}C(e)}{\sqrt{1-e^{2}}}.$
Due to $\frac{G(e)}{\sqrt{1-e^{2}}}>1$ for all $e\in(0,1)$, condition (c) holds. ∎
6. Appendix. The Maslov-type index for symplectic paths
In this section, we briefly review the index theory for symplectic path. The detail could be found in [15]. Let $(\mathbb{R}^{2n},\omega)$ be the standard symplectic vector space with coordinates $(x_{1},\cdots,x_{n},y_{1},\cdots,y_{n})$, then $\omega=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$. Let $J=\begin{pmatrix}0&-I_{n}\\ I_{n}&0\end{pmatrix}$, where $I_{n}$ is the identity matrix on $\mathbb{R}^{n}$. The symplectic group $\text{Sp}(2n)$ is defined by
$\text{Sp}(2n)=\{M\in GL(2n,\mathbb{R})\big{|}M^{T}JM=J\},$
with topology induced from $\mathbb{R}^{4n^{2}}$. For $\tau>0$, we define the symplectic paths set
$\mathcal{P}_{\tau}(2n)=\{\gamma\in C([0,\tau],\text{Sp}(2n))\big{|}\gamma(0)=I_{2n}\}.$
For any $\omega\in\mathbb{U}$ (the unite circle in $\mathbb{C}$), the following $\omega$-degenerate hypersurface of codimension one in $\text{Sp}(2n)$ is defined by
$\text{Sp}(2n)_{\omega}^{0}=\{M\in\text{Sp}(2n)\big{|}\det(M-\omega I_{2n})=0\}.$
Moreover, the $\omega$-regular set of $\text{Sp}(2n)$ is defined by $\text{Sp}(2n)_{\omega}^{*}=\text{Sp}(2n)\setminus\text{Sp}(2n)_{\omega}^{0}$. For any two continuous path $\xi$ and $\eta:[0,\tau]\rightarrow\text{Sp}(2n)$ with $\xi(\tau)=\eta(0)$, we define their concatenation as:
$\eta*\xi(t)=\begin{cases}\xi(2t),&\text{ if }0\leq t\leq\tau/2;\\ \eta(2t-\tau),&\text{ if }\tau/2\leq t\leq\tau.\end{cases}$
Given any two $2m_{k}\times 2m_{k}$ matrices of square block form $M_{k}=\begin{pmatrix}A_{k}&B_{k}\\ C_{k}&D_{k}\end{pmatrix}$ with $k=1,2$, the $\diamond$-product of $M_{1}$ and $M_{2}$ defined by the following $2(m_{1}+m_{2})\times 2(m_{1}+m_{2})$ matrix $M_{1}\diamond M_{2}$:
$M_{1}\diamond M_{2}=\begin{pmatrix}A_{1}&0&B_{1}&0\\ 0&A_{2}&0&B_{2}\\ C_{1}&0&D_{1}&0\\ 0&C_{2}&0&D_{2}\end{pmatrix}$.
We define a special continuous symplectic path $\xi_{m}\subset\text{Sp}(2n)$ by
$\xi_{n}(t)=\begin{pmatrix}2-\frac{t}{\tau}&0\\ 0&\left(2-\frac{t}{\tau}\right)^{-1}\end{pmatrix}^{\diamond n}\text{ for }0\leq t\leq\tau.$
For $M\in\text{Sp}(2n)_{\omega}^{0}$, we define a co-orientation of $\text{Sp}(2n)_{\omega}^{0}$ at $M$ by the positive direction $\frac{\mathrm{d}}{\mathrm{d}t}Me^{tJ}\big{|}_{t=0}$ of the path $Me^{tJ}$ with $|t|$ sufficiently small. Now we will give the definition of $\omega$-index.
Definition 6.1.
For $\omega\in\mathbb{U}$, $\gamma\in\mathcal{P}_{\tau}(2n)$, the $\omega$-index of $\gamma$ is defined as
$i_{\omega}(\gamma)=\left[e^{-\varepsilon J}\gamma*\xi_{n}:\text{Sp}(2n)_{\omega}^{0}\right],$
for $\varepsilon>0$ small enough, where $\cdot$$\cdot$] is the intersection number. The $\omega$-nullity of $\gamma$ is defined as
$\nu_{\omega}(\gamma)=\dim_{\mathbb{C}}\ker_{\mathbb{C}}(\gamma(\tau)-\omega I_{2n}).$
Particularly, if $\gamma(\tau)\in\text{Sp}(2n)_{\omega}^{*}$, we can take $\varepsilon=0$.
For any two path $\gamma_{j}\in\mathcal{P}_{\tau}(2n_{j})$ with $j=1,2$, let $\gamma_{1}\diamond\gamma_{2}(t)=\gamma_{1}(t)\diamond\gamma_{2}(t)$ for all $t\in[0,\tau]$. We can easily obtain
$i_{\omega}(\gamma_{1}\diamond\gamma_{2})=i_{\omega}(\gamma_{1})+i_{\omega}(\gamma_{2}),\quad\forall\omega\in\mathbb{U}.$
For any symplectic path $\gamma\in\mathcal{P}_{\tau}(2n)$ and $m\in\mathbb{Z}^{+}$, we define the $m$-th iteration $\gamma^{m}:[0,m\tau]\rightarrow\text{Sp}(2n)$ by
$\gamma^{m}(t)=\gamma(t-j\tau)\gamma(\tau)^{j},\quad\text{ for }j\tau\leq t\leq(j+1)\tau,j=0,1,\cdots m-1.$
Definition 6.2.
The mean index of $\gamma$ is defined as
$\hat{i}(\gamma)=\lim\limits_{m\rightarrow+\infty}\frac{i_{1}(\gamma^{m})}{m}.$
Let $x(t)$ be a periodic solution of Hamiltonian system $\dot{x}=J\nabla H(x)$ with period $\tau$, then the fundamental solution of the linearized equation
$\dot{y}=JD^{2}H(x(t))y$
is a symplectic path starting form $I_{2n}$, denote it by $\gamma(t)$. Then we can define the $\omega$-index and mean index for $x(t)$.
Definition 6.3.
The $\omega$-index and $\omega$-nullity of $x(t)$ is defined as
$i_{\omega}(x)=i_{\omega}(\gamma),\quad\nu_{\omega}(x)=\nu_{\omega}(\gamma).$
The mean index of $x(t)$ is defined as
$\hat{i}(x)=\hat{i}(\gamma).$
For mechanical systems with $\gamma\in\mathcal{P}_{\tau}(2)$, the analysis is simpler than the general case. Let $\sigma(t)=\rho(t)e^{i\vartheta(t)}$ and $\sigma^{-1}(t)$ denote the eigenvalues of $\gamma(t)$, where $\vartheta(t)$ is a continuous non-decreasing function. The mean index of $\gamma$ is given by
$\hat{i}(\gamma)=\frac{\vartheta(\tau)-\vartheta(0)}{\pi}.$